### v2

# Set-Valued Koopman Theory for Control Systems

Abstract
In this paper, we introduce a new notion of Koopman operator which faithfully encodes the dynamics of controlled systems by leveraging the grammar of set-valued analysis. In this context, we propose meaningful generalisations of the Liouville and Perron-Frobenius operators, and show that they respectively coincide with proper set-valued analogues of the infinitesimal generator and dual operator of the Koopman semigroup. We also give meaning to the spectra of these set-valued maps and prove an adapted version of the classical spectral mapping theorem relating the eigenvalues of a semigroup and those of its generator. In essence, these results provide theoretical justifications for existing approaches in the Koopman communities which consist in studying control systems by bundling together the Liouville operators associated with different input parameters.
Keywords : Koopman Operators, Control Systems, Differential Inclusions, Set-Valued Analysis.
MSC2020 Subject Classification : 28B20, 34A60, 47N20, 93C15.
1 Introduction
The Koopman operator was introduced in functional analysis as a way to provide an equivalent representation of nonlinear time-evolutions – or more generally of flow maps – in terms of infinite-dimensional linear operators, and is now a well-established framework for the analysis of dynamical systems. Originating from the seminal papers of Koopman [29] and Koopman and von Neumann [28] in the early 1930s, it has enjoyed a renewed interest pioneered by the works of Mezić and Banaszuk [37] and Mezić [36]. Heuristically, given a one-parameter semigroup $(\Phi_{t})_{t\geq 0}$ representing e.g. the integral curves of a dynamical system, the Koopman operators are linear transforms whose action is given by
$\pazocal{K}_{t}(\varphi):=\varphi\circ\Phi_{t}$ (1.1)
for each (typically real or complex valued) function $\varphi$ belonging to some relevant subspace of function observables, denoted by $\pazocal{X}$. Besides its appealing theoretical properties, which allows for the use of the whole corpus of spectral theory available for linear operators to understand nonlinear systems, the Koopman operator has served as a foundation for state of the art numerical methods in computational mathematics, most notably the (Extended) Dynamic Mode Decomposition [46, 54].
More recently, the Koopman operator framework has been used for systems with external inputs, first regarded as exogenous disturbances [41] that one cannot manipulate, and later, starting with the work [30], as controls that one may tune in order to achieve specific objectives. When used within a model predictive control framework, the approach proposed in the latter work took the form of a convex optimization problem, contrarily to traditional nonlinear model predictive schemes. This very appealing feature has led to a number of follow-up works aiming at improving the practical aspects of the method, see e.g. [14, 31, 39, 40, 48], as well as a number of applications including soft robotics [25], power grid stabilisation [32], fluids [3, 40] or quantum control [24]. This list is by no means exhaustive, and we point the interested reader to the surveys [7, 12] for more references. However, contrary to the abundance of methodological advances and applications of the Koopman framework, a sound theoretical footing for the Koopman operator associated with controlled dynamics is still missing at present. Indeed, the work [30] defined the Koopman operator with control on the so-called tensor-product system but did not leverage this definition for theoretical analysis. In [41], the authors considered the Koopman operator with one fixed value of the control input or with control signals subject to a dynamical evolution (e.g., determined by a feedback law), again without providing theoretical insights, while the subsequent work [40] focused on the Koopman operators associated with each individual control input and interpolated between them for control design purposes. The work [43] likewise considered control-affine systems and defined the control Koopman operator as the collection of the Koopman operators corresponding to the drift and each of the control vector fields.
In this article, we develop a theoretical framework for Koopman operators in the presence of controls, based on the theory of set-valued analysis. Starting from the seminal works of Filippov [20] and Wazewski [53] at the turn of the 1960s, it has been known that rephrasing control problems in terms of differential inclusions provided key insights on the optimal sets of assumptions needed to establish positive results on controlled dynamics, while also bringing powerful tools from geometric and nonsmooth analysis to investigate such systems. Since then, the methods of set-valued analysis have been successfully applied to a large breadth of control problems throughout the past decades, ranging from the well-posedness of constrained dynamical systems both in the classical [6] and hybrid [23] settings, to Pontryagin [16, 51] and Hamilton-Jacobi [21, 22] optimality conditions as well as Lyapunov stability methods [49]. This list of references is far from complete, and we point the interested reader to the reference monographs of Aubin, Cellina, Clarke, Frankowska and Vinter [4, 5, 15, 50] for further details. Based on these observations, we propose a comprehensive adaptation of the main concepts of Koopman theory to general time-invariant control systems of the form
$\dot{x}(t)=f(x(t),u(t)),$
where $f:\mathbb{R}^{d}\times U\to\mathbb{R}^{d}$ is locally Lipschitz and sublinear in $x\in\mathbb{R}^{d}$ as well as continuous in $u\in U$. In this context, one can associate to each admissible control signal $u(\cdot)\in\pazocal{U}$ a unique family of flow maps $(\Phi_{(\tau,t)}^{u})_{\tau,t\geq 0}\subset C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ satisfying
$\Phi_{(\tau,t)}^{u}(x)=x+\int_{\tau}^{t}f\Big{(}\Phi_{(\tau,s)}^{u}(x),u(s)\Big{)}\textnormal{d}s$
for all times $\tau,t\geq 0$ and every $x\in\mathbb{R}^{d}$. This leads us to defining the set-valued Koopman operators as the collection of evaluations of a given observable $\varphi\in\pazocal{X}$ along all possible controlled flows, that is
$\pazocal{K}_{(\tau,t)}(\varphi):=\Big{\{}\varphi\circ\Phi_{(\tau,t)}^{u}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u(\cdot)\in\pazocal{U}\Big{\}}.$
In other words, $\pazocal{K}_{(\tau,t)}(\varphi)$ is the set of all observables which are reachable using admissible control trajectories starting from $\varphi\in\pazocal{X}$. We discuss some of the fundamental topological properties of these objects in Section 3, and provide a general representation formula for arbitrary time-dependent Koopman observables $(\tau,t)\in[0,T]\times[0,T]\mapsto\psi_{(\tau,t)}\in\pazocal{K}_{(\tau,t)}(\varphi)$, involving measurable families of control signals. It should be noted that while autonomous controlled systems can be described without loss of generality by a one-parameter semigroup, up to a time reparametrisation, we chose to define the Koopman operators as a two-parameter family in the spirit of [34]. The main incentive for doing so lies in the fact that the dynamics of Koopman observables depends on the starting time of the controlled flows, as shown in the next paragraph and more thoroughly in Section 4.1 below.
[…]
Figure 1.1: The classical Koopman operator associates to an observable $\varphi:\mathbb{R}^{d}\to\mathbb{C}$ the measurement $\varphi\circ\Phi_{t}(x)$ along a single flow $\Phi_{t}(x)$ evaluated from every possible initial data $x\in\mathbb{R}^{d}$ at time $t\in[0,T]$ (left), whereas the set-valued Koopman operator outputs the whole reachable set $\{\varphi\circ\Phi_{(0,t)}^{u}(x)\}_{u(\cdot)\in\pazocal{U}}$ of observables along the collection of controlled flows $\{\Phi_{(0,t)}^{u}(x)\}_{u(\cdot)\in\pazocal{U}}$ starting from any $x\in\mathbb{R}^{d}$ and evaluated at time $t\in[0,T]$ (right).
In Section 4, which is the core of the manuscript, we put forth relevant set-valued counterparts of the Liouville and Perron-Frobenius operators. These objects are known to play a pivotal role in Koopman theory as well as many of its applications, the former by being the infinitesimal generator of the Koopman semigroup and therefore encoding the dynamics of Koopman observables, and the latter by being the topological dual of the Koopman operators, thus providing alternative angles to address dynamical systems problems. In Section 4.1, we define the set-valued Liouville operators as
$\pazocal{L}(\varphi)=\Big{\{}\nabla_{x}\varphi\cdot f_{u}\leavevmode\nobreak\ \,\mathrm{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}$
for each continuously differentiable observable $\varphi\in\pazocal{D}$ and where $f_{u}\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ is given by $f_{u}(x):=f(x,u)$ for all $(x,u)\in\mathbb{R}^{d}\times U$. We then prove that, when the set of controlled vector fields is convex, the Liouville operator is the infinitesimal generator of the Koopman semigroup, in the sense that
$\underset{\leavevmode\nobreak\ t\to\tau}{\textnormal{Lim}}\frac{\pazocal{K}_{(\tau,t)}(\varphi)-\varphi}{t-\tau}=\pazocal{L}(\varphi),$
where limit is understood in the sense of Kuratowski-Painlevé. In addition, when $f$ is continuously differentiable in $x$ for every $u\in U$, we prove that time-dependent Koopman observables given by $(\tau,t)\in[0,T]\times[0,T]\mapsto\psi_{(\tau,t)}:=\varphi\circ\Phi_{(\tau,t)}^{u}\in\pazocal{D}$ for some fixed control signal $u(\cdot)\in\pazocal{U}$ coincide exactly with the strong solutions of the functional differential inclusion
[…]
in the space of observables $\pazocal{X}$. In Section 4.2, we shift our focus to the investigation of duality results for the set-valued Koopman operators. In this context, we propose the following definition for the set-valued Perron-Frobenius operators
$\pazocal{P}_{(\tau,t)}(\mu):=\Big{\{}\Phi_{(\tau,t)\sharp\,}^{u}\mu\leavevmode\nobreak\ \,\text{s.t.}\leavevmode\nobreak\ u(\cdot)\in\pazocal{U}\Big{\}},$
where $\sharp$” stands for the usual image measure operation, and show that the weak- convex hull of the latter coincide with the set-valued adjoints of the Koopman operators, defined in the sense of Ioffe [26]. We also leverage this notion of duality to prove that, when the dynamics is convex, the infinitesimal generator of the Perron-Frobenius semigroup is the adjoint of the Liouville operator, namely
$\underset{\leavevmode\nobreak\ t\to\tau}{\textnormal{Lim}}\frac{\pazocal{P}_{(\tau,t)}(\mu)-\mu}{t-\tau}=\pazocal{L}^{*}(\mu)$
in the weak- topology of $\pazocal{X}^{*}$. To our surprise, we discovered that the underlying functional inclusion
[…]
coincided – at least on a formal level – with the notion of continuity inclusion introduced by the first author and Frankowska in [8, 9] in the context of mean-field control. Lastly, in Section 4.3, we prove a set-valued version of the usual spectral mapping theorem (see e.g. [19, Chapter IV - Theorem 3.7]), which relates the point spectra of the Liouville operator with that of the Koopman semigroup. Therein, we show in particular that
$e^{(t-\tau)\sigma_{p}(\pazocal{L})}\subset\sigma_{p}(\pazocal{K}_{(\tau,t)}),$
which provides grounding to preexisting works such as [39, 40] in which spectral properties of Koopman operators for control systems are investigated by exponentiating the spectra of the collection of Liouville operators associated with a finite collection of admissible control values.
The manuscript is organised as follows. In Section 2 we start by exposing preliminary notions of functional and set-valued analysis as well as control theory. In Section 3, we define the set-valued Koopman operators and study their main properties. We then move to the investigation of the set-valued Liouville and Perron-Frobenius operators in Section 4, wherein we show that the latter are respectively the infinitesimal generator and related to the dual of the Koopman semigroups. We then close this section with the statement of a set-valued counterpart of spectral mapping theorem, and provide the proof of a technical regularity result for controlled flows in Appendix A.
2 Preliminaries
In this first section, we collect preliminary material on integration theory, set-valued analysis and controlled systems, for which we point the reader to the monographs [18], [5] and [15] respectively.
2.1 Functional analysis and integration
In what follows, we denote by $(\Omega,\pazocal{A},\mathfrak{m})$ a $\sigma$-finite measure space, and recall following [5, Section 8.1] that a map $f:\Omega\to\pazocal{X}$ valued in a complete separable metric space $(\pazocal{X},\textnormal{{d}}_{\pazocal{X}}(\cdot,\cdot))$ is $\mathfrak{m}$-measurable if
$f^{-1}(\pazocal{O}):=\Big{\{}\omega\in\Omega\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ f(\omega)\in\pazocal{O}\Big{\}}\subset\Omega$
is $\mathfrak{m}$-measurable for each open set $\pazocal{O}\subset\pazocal{X}$. Throughout the manuscript, we shall denote by $\mathscr{L}^{d}$ the standard Lebesgue measure defined over $(\mathbb{R}^{d},\pazocal{B}(\mathbb{R}^{d}))$. Below, we recall the concept of integrability in the sense of Bochner for maps valued in separable Banach spaces, see for instance [18, Chapter II].
Definition 2.1 (Bochner integrable maps).
An $\mathfrak{m}$-measurable map $f:\Omega\to\pazocal{X}$ valued in a separable Banach space $(\pazocal{X},\|\cdot\|_{\pazocal{X}})$ is said to be Bochner integrable if it satisfies
$\int_{\Omega}\parallel f(\omega)\parallel_{\pazocal{X}}\textnormal{d}\mathfrak{m}(\omega)<+\infty.$
The collection of all such maps is a separable Banach space that we denote by $L^{1}(\Omega,\pazocal{X})$.
Similarly to the classical Lebesgue integral, the Bochner integral of a map against a measure with unit mass defines a mapping whose range is convex, as shown by the following proposition.
Proposition 2.2 (Convexity of the range of the Bochner integral).
Let $(\pazocal{X},\parallel\cdot\parallel_{\pazocal{X}})$ be a separable Banach space, $K\subset\pazocal{X}$ be a closed set and $f(\cdot)\in L^{1}(\Omega,\pazocal{X})$ be such that $f(\omega)\in K$ for $\mathfrak{m}$-almost every $\omega\in\Omega$. Then, if $\mathfrak{m}(\Omega)=1$, it holds that
$\int_{\Omega}f(\omega)\textnormal{d}\mathfrak{m}(\omega)\in\overline{\textnormal{co}}\hskip 1.42271ptK.$
Proof.
This result follows from a straightforward combination of the Hahn-Banach separation theorem (see e.g. [11, Theorem 1.7]) together with the linearity and homogeneity of the Bochner integral. ∎
Given a locally convex topological vector space $\pazocal{X}$, we denote by $\pazocal{X}^{*}$ its topological dual, i.e. the collection of all bounded complex-valued linear functionals over $\pazocal{X}$, and write $\langle\cdot,\cdot\rangle_{\pazocal{X}}$ for the underlying duality pairing, see e.g. [45, Part 1 – Chapter 1]. Throughout the paper, we will systematically assume that $\pazocal{X}^{*}$ is endowed with the corresponding weak- topology, see e.g. [45, Part 1 – Chapter 3]. In the next proposition, we recall a fine compactness criterion for the weak topology of $L^{1}(\Omega,\pazocal{X})$, which is excerpted from [17, Corollary 2.6].
Proposition 2.3 (A weak $L^{1}$-compactness criterion for the Bochner integral).
Let $(\pazocal{X},\parallel\cdot\parallel_{\pazocal{X}})$ be a separable Banach space and $(f_{n}(\cdot))\subset L^{1}(\Omega,\pazocal{X})$. Suppose that there exists a map $k(\cdot)\in L^{1}(\Omega,\mathbb{R}_{+})$ along with a convex and compact set $K\subset\pazocal{X}$ such that
$\parallel f_{n}(\omega)\parallel_{\pazocal{X}}\,\leq k(\omega)\qquad\text{and}\qquad f_{n}(\omega)\in K$
for $\mathfrak{m}$-almost every $\omega\in\Omega$. Then, there exists a subsequence $(f_{n_{k}}(\cdot))$ that converges weakly to some $f(\cdot)\in L^{1}(\Omega,\pazocal{X})$ satisfying $f(\omega)\in K$ for $\mathfrak{m}$-almost every $\omega\in\Omega$. In particular, it holds that
$\int_{\Omega}\big{\langle}\bm{\nu}(\omega),f(\omega)-f_{n_{k}}(\omega)\big{\rangle}_{\pazocal{X}}\,\textnormal{d}\mathfrak{m}(\omega)\leavevmode\nobreak\ \underset{k\to+\infty}{\longrightarrow}\leavevmode\nobreak\ 0$
for each $\bm{\nu}:\Omega\to\pazocal{X}^{*}$ such that $\omega\in\Omega\mapsto\langle\bm{\nu}^{*}(\omega),f\,\rangle_{\pazocal{X}}\in\mathbb{R}$ is $\mathfrak{m}$-measurable for every $f\in\pazocal{X}$ with $\textnormal{ess sup}_{\omega\in\Omega}\parallel\bm{\nu}^{*}(\omega)\parallel_{\pazocal{X}^{*}}<+\infty$.
Given some $k\in\{0,1\}$, we denote by $(C^{k}_{c}(\mathbb{R}^{d},\mathbb{C}),\left\|\cdot\right\|_{C^{k}(\mathbb{R}^{d},\mathbb{C})})$ the separable normed space of $k$-times continuously differentiable maps with compact support, endowed with the relevant supremum norm. By Riesz’s representation theorem (see e.g. [44, Theorem 6.19]), one has that
$C_{c}^{0}(\mathbb{R}^{d},\mathbb{C})^{*}\simeq\mathscr{M}(\mathbb{R}^{d},\mathbb{C})$
where $\mathscr{M}(\mathbb{R}^{d},\mathbb{C})$ is the vector space of finite complex-valued Radon measures. In this context, given an element $\mu\in\mathscr{M}(\mathbb{R}^{d},\mathbb{C})$ and some $f\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$, we define the so-called divergence distribution $\textnormal{div}_{x}(f\mu)\in C^{1}_{c}(\mathbb{R}^{d},\mathbb{C})^{*}$ via its duality action as
$\langle\textnormal{div}_{x}(f\mu),\zeta\rangle_{C^{1}_{c}(\mathbb{R}^{d},\mathbb{C})}:=-\int_{\mathbb{R}^{d}}\nabla_{x}\zeta(x)\cdot f(x)\textnormal{d}\mu(x)$ (2.1)
for all $\zeta\in C^{1}_{c}(\mathbb{R}^{d},\mathbb{C})$. We also recall that the image – or pushforward – of a finite Radon measure $\mu\in\mathscr{M}(\mathbb{R}^{d},\mathbb{C})$ through a Borel map $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the unique measure $f_{\sharp}\mu\in\mathscr{M}(\mathbb{R}^{d},\mathbb{C})$ which satisfies
$f_{\sharp}\mu(B):=\mu(f^{-1}(B))$
for each Borel set $B\subset\mathbb{R}^{d}$. Moreover, the latter is characterised by the change of variable formula
$\int_{\mathbb{R}^{d}}\varphi(x)\textnormal{d}(f_{\sharp}\mu)(x)=\int_{\mathbb{R}^{d}}\varphi\circ f(x)\textnormal{d}\mu(x)$ (2.2)
which holds for every $\varphi\in C^{0}_{c}(\mathbb{R}^{d},\mathbb{C})$. In the next definition, we recollect the notion of weak- convergence for finite Radon measures (see e.g. [2, Definition 1.58]).
Definition 2.4 (Weak- convergence of measures).
We say that a sequence $(\mu_{n})\subset\mathscr{M}(\mathbb{R}^{d},\mathbb{C})$ converges towards some $\mu\in\mathscr{M}(\mathbb{R}^{d},\mathbb{C})$ for the weak- topology provided that
$\int_{\mathbb{R}^{d}}\zeta(x)\textnormal{d}\mu_{n}(x)\leavevmode\nobreak\ \underset{n\to+\infty}{\longrightarrow}\leavevmode\nobreak\ \int_{\mathbb{R}^{d}}\zeta(x)\textnormal{d}\mu(x)$ (2.3)
for each $\zeta\in C^{0}_{c}(\mathbb{R}^{d},\mathbb{C})$.
In what ensues, we will also consider elements of the space $C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ of continuous functions defined over the whole space. In that case, the adequate topology to consider is that of local uniform convergence, whose definition is recalled below. Therein and throughout the article, we write $B(0,R)\subset\mathbb{R}^{d}$ for the closed Euclidean ball of radius $R>0$ centered at the origin.
Definition 2.5 (Topology of uniform convergence on compact sets).
A sequence of maps $(f_{n})\subset C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ converges uniformly on compact sets – or locally uniformly – towards $f\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ if
[…]
for each compact set $K\subset\mathbb{R}^{d}$. The underlying topology is induced by the translation invariant metric
$\textnormal{{d}}_{cc}(f,g):=\sum_{k=1}^{+\infty}2^{-k}\min\Big{\{}1\,,\,\left\|f-g\right\|_{C^{0}(B(0,k),\mathbb{R}^{d})}\Big{\}},$
which is defined for each $f,g\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$.
Following the general results from [27, Chapter 7, Theorems 12 and 13] and [52, Theorem 6], it can be shown $(C^{0}(\mathbb{R}^{d},\mathbb{R}^{d}),\textnormal{{d}}_{cc}(\cdot,\cdot))$ is a complete separable metric space. More classically, it is well known that $(C^{0}(K,\mathbb{R}^{d}),\left\|\cdot\right\|_{C^{0}(K,\mathbb{R}^{d})})$ is a separable Banach space for each compact set $K\subset\mathbb{R}^{d}$.
2.2 Set-valued analysis
In this section, we recall preliminary material pertaining to set-valued analysis, for which we point the reader to the reference monograph [5]. In what follows unless further specifications are given, we suppose that $\pazocal{X}$ and $\pazocal{Y}$ are locally convex topological vector spaces.
Definition 2.6 (Set-valued maps).
We say that $\pazocal{F}:\pazocal{X}\rightrightarrows\pazocal{Y}$ is a set-valued map – or multifunction – if $\pazocal{F}(x)\subset\pazocal{Y}$ for each $x\in\pazocal{X}$. Its domain is then defined by
$\textnormal{dom}(\pazocal{F}):=\Big{\{}x\in\pazocal{X}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ \pazocal{F}(x)\neq\emptyset\Big{\}}.$
Below, we recall classical regularity notions for set-valued mappings, starting with the standard concepts of continuity and Lipschitz continuity, and proceed with that of measurability. Therein, we suppose that $(\pazocal{X},\textnormal{{d}}_{\pazocal{X}}(\cdot,\cdot))$ and $(\pazocal{Y},\textnormal{{d}}_{\pazocal{Y}}(\cdot,\cdot))$ are complete separable metric spaces, write $\mathbb{B}_{\pazocal{X}}(x,r)$ for the closed ball of radius $r>0$ centered at some $x\in\pazocal{X}$ and $\mathbb{B}_{\pazocal{X}}(\pazocal{Q},r):=\{y\in\pazocal{X}\leavevmode\nobreak\ \,\text{s.t.}\leavevmode\nobreak\ \textnormal{{d}}_{\pazocal{X}}(x,y)\leq r\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for some $x\in\pazocal{Q}$}\}$ for the $r$-neighbourhood of a set $\pazocal{Q}\subset\pazocal{X}$.
Definition 2.7 (Continuity of set-valued maps).
A multifunction $\pazocal{F}:\pazocal{X}\rightrightarrows\pazocal{Y}$ is said to be continuous at $x\in\textnormal{dom}(\pazocal{F})$ if both the following conditions hold.
  1. $(i)$
    $\pazocal{F}$ is lower-semicontinuous at $x$, i.e. for any $\varepsilon>0$ and all $y\in\pazocal{F}(x)$, there exists $\delta>0$ such that
    $\pazocal{F}(x^{\prime})\cap\mathbb{B}_{\pazocal{Y}}(y,\varepsilon)\neq\emptyset$
    for each $x^{\prime}\in\mathbb{B}_{\pazocal{X}}(x,\delta)$.
  2. $(ii)$
    $\pazocal{F}$ is upper-semicontinuous at $x$, i.e. for any $\varepsilon>0$, there exists $\delta>0$ such that
    $\pazocal{F}(x^{\prime})\subset\mathbb{B}_{\pazocal{Y}}(\pazocal{F}(x),\varepsilon)$
    for each $x^{\prime}\in\mathbb{B}_{\pazocal{X}}(x,\delta)$.
Definition 2.8 (Lipschitz continuity of set-valued maps).
A multifunction $\pazocal{F}:\pazocal{X}\rightrightarrows\pazocal{Y}$ is Lipschitz continuous with constant $L>0$ provided that
$\pazocal{F}(x^{\prime})\subset\mathbb{B}_{\pazocal{Y}}\Big{(}\pazocal{F}(x),L\,\textnormal{{d}}_{\pazocal{X}}(x,x^{\prime})\Big{)}$ (2.4)
for all $x,x^{\prime}\in\textnormal{dom}(\pazocal{F})$. In particular, $\pazocal{F}$ is continuous at every point $x\in\textnormal{dom}(\pazocal{F})$.
Definition 2.9 (Measurable set-valued maps and selections).
A multifunction $\pazocal{F}:\Omega\rightrightarrows\pazocal{X}$ is said to be $\mathfrak{m}$-measurable – or more simply measurable – provided that for any open set $\pazocal{O}\subset\pazocal{X}$, the preimage
$\pazocal{F}^{-1}(\pazocal{O}):=\Big{\{}\omega\in\Omega\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ \pazocal{F}(\omega)\cap\pazocal{O}\neq\emptyset\Big{\}}$
is $\mathfrak{m}$-measurable. Moreover, a measurable map $f:\Omega\to\pazocal{X}$ is called a measurable selection of $\pazocal{F}:\Omega\rightrightarrows\pazocal{X}$ provided that $f(\omega)\in\pazocal{F}(\omega)$ for $\mathfrak{m}$-almost every $\omega\in\Omega$.
One may note that whenever $\pazocal{F}(\omega)=\{f(\omega)\}$ is single-valued, this definition amounts to requiring that $f:\Omega\to\pazocal{X}$ be Lebesgue measurable in the sense introduced above, since then
$\pazocal{F}^{-1}(\pazocal{O})=\Big{\{}\omega\in\Omega\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ \{f(\omega)\}\cap\pazocal{O}\neq\emptyset\Big{\}}=f^{-1}(\pazocal{O}).$
In the following theorem, we recall a variant of the instrumental Filippov selection principle, whose statement may be found e.g. in [5, Theorem 8.2.10].
Theorem 2.10 (Filippov’s selection theorem).
Let $\pazocal{F}:\Omega\rightrightarrows\pazocal{X}$ be an $\mathfrak{m}$-measurable set-valued map with nonempty closed images and $\Psi:\Omega\times\pazocal{X}\to\pazocal{Y}$ be $\mathfrak{m}$-measurable in $\omega\in\Omega$ and continuous in $x\in\pazocal{X}$. Then for every $\mathfrak{m}$-measurable map $\psi:\Omega\to\pazocal{Y}$ satisfying
$\psi(\omega)\in\Psi(\omega,\pazocal{F}(\omega)):=\Big{\{}\Psi(\omega,f)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ f\in\pazocal{F}(\omega)\Big{\}},$
there exists a measurable selection $\omega\in\Omega\mapsto f(\omega)\in\pazocal{F}(\omega)$ such that $\psi(\omega)=\Psi(\omega,f(\omega))$ for $\mathfrak{m}$-almost every $\omega\in\Omega$.
In what follows, we recollect the definitions of standard limit concepts for sequences of sets. Therein, we shall use the notation
$\textnormal{dist}_{\pazocal{X}}(x\,;K):=\inf_{y\in K}\textnormal{{d}}_{\pazocal{X}}(x,y)$
for the distance between a point $x\in\pazocal{X}$ and a set $K\subset\pazocal{X}$.
Definition 2.11 (Kuratowski-Painlevé limit of sequences of sets).
Given a sequence of sets $(K_{n})$, we define its lower limit as
$\underset{\leavevmode\nobreak\ n\to+\infty}{\textnormal{Liminf}}\,K_{n}:=\bigg{\{}x\in\pazocal{X}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ \lim_{n\to+\infty}\,\textnormal{dist}_{\pazocal{X}}(x,K_{n})=0\bigg{\}},$
as well as its upper limit by
$\underset{\leavevmode\nobreak\ n\to+\infty}{\textnormal{Limsup}}\,K_{n}:=\bigg{\{}x\in\pazocal{X}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ \liminf_{n\to+\infty}\,\textnormal{dist}_{\pazocal{X}}(x,K_{n})=0\bigg{\}}.$
In this context, we say that $K\subset\pazocal{X}$ is the Kuratowski-Painlevé limit of the sequence $(K_{n})$ provided
$\underset{\leavevmode\nobreak\ n\to+\infty}{\textnormal{Lim}}\,K_{n}:=\underset{\leavevmode\nobreak\ n\to+\infty}{\textnormal{Liminf}}\,K_{n}=\underset{\leavevmode\nobreak\ n\to+\infty}{\textnormal{Limsup}}\,K_{n}=K.$
[…]
Figure 2.1: Illustration of the concepts of lower and upper limits for sequences of sets. The lower and upper limits of the sequence with general term $K_{n}:=\{\tfrac{1}{n}\}\times[0,(-1)^{n}]$ (left) are respectively $\{(0,0)\}$ and $\{0\}\times[-1,1]$, whereas those of the sequence $K_{n}:=\{\tfrac{1}{n}\}\times[-\tfrac{1}{n},\tfrac{1}{n}]$ (right) coincide and are equal to the singleton $\{(0,0)\}$.
We close this section by recollecting the definition of closed processes and fans between locally convex topological vector spaces. The latter were introduced respectively by Rockafellar at the end of the 1960s, see e.g. [42, Chapter 9], and Ioffe in the 1980’s, see e.g. [26], with the goal to adapt the main and most appealing properties of closed and linear operators to the set-valued setting.
Definition 2.12 (Closed processes and fans).
A set-valued map $\pazocal{F}:\pazocal{X}\rightrightarrows\pazocal{Y}$ is called a closed process if its graph $\textnormal{Graph}(\pazocal{F})\subset\pazocal{X}\times\pazocal{Y}$ is a closed cone. A closed process $\pazocal{F}:\pazocal{X}\rightrightarrows\pazocal{Y}$ with convex images is called a fan provided that
$0\in\pazocal{F}(0)\qquad\text{and}\qquad\pazocal{F}(x_{1}+x_{2})\subset\pazocal{F}(x_{1})+\pazocal{F}(x_{2})$
for each $x_{1},x_{2}\in\textnormal{dom}(\pazocal{F})$, namely if it is homogeneous and subadditive.
Example 2.13 (Collections of linear operators as closed processes and fans).
To see why these objects can indeed be seen as generalisations of bounded linear operators, suppose that $\pazocal{F}(x):=\{Ax\}$ for each $x\in\pazocal{X}$ and some bounded linear map $A:\pazocal{X}\to\pazocal{Y}$. Then, one may easily check that
$\textnormal{Graph}(\pazocal{F})=\big{\{}(x,Ax)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ x\in\pazocal{X}\big{\}}=\textnormal{Graph}(A)$
is an additive closed cone and that the images of $\pazocal{F}:\pazocal{X}\rightrightarrows\pazocal{Y}$ are convex, hence the latter defines a fan. In the case in which $A:\textnormal{dom}(A)\to\pazocal{Y}$ is a linear operator defined over some domain $\textnormal{dom}(A)\subset\pazocal{X}$, the mapping $\pazocal{F}:\pazocal{X}\rightrightarrows\pazocal{Y}$ is also a fan. More interestingly, if $\pazocal{A}$ is a convex and closed collection of linear operators from $\pazocal{X}$ to $\pazocal{Y}$, then $\pazocal{F}(x):=\{Ax\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ A\in\pazocal{A}\,\}$ given for each $x\in\pazocal{X}$ defines a fan.
$\pazocal{X}$$\pazocal{Y}$$\pazocal{X}$$\pazocal{Y}$
Figure 2.2: Illustration of the concepts of closed processes and fans. The set-valued mapping given by $\pazocal{F}(x):=[-x,x]$ (left) is clearly homogeneous and convex-valued as well as subadditive, and thus defines a fan. On the other hand, the multifunction given by $\pazocal{F}(x)=[-x,-3x/4]\cup[-x/2,x/2]\cup[3x/4,x]$ (right) is a homogeneous closed process, but has non-convex images and is therefore not a fan.
2.3 Controlled dynamical systems
In what follows, we recollect elementary facts pertaining to controlled Cauchy problems of the form
[…]
defined over some finite time interval $[0,T]$ with horizon $T>0$, and taking a prescribed value $x\in\mathbb{R}^{d}$ at some time $\tau\in[0,T]$. Therein and in what follows, the set of admissible controls is given by
[…]
for some compact metric space $(U,\textnormal{{d}}_{U}(\cdot,\cdot))$, where $\mathscr{L}^{1}$ stands for the standard Lebesgue measure over $[0,T]$. Throughout the paper, we will assume that the following assumptions hold.
Hypotheses (H).
  1. $(i)$
    The map $f:(x,u)\in\mathbb{R}^{d}\times U\to\mathbb{R}^{d}$ is continuous, and there exists a constant $m>0$ such that
    $|f(x,u)|\leq m\big{(}1+|x|\big{)}$
    for all $(x,u)\in\mathbb{R}^{d}\times U$.
  2. $(ii)$
    For each compact set $K\subset\mathbb{R}^{d}$, there exists a constant $\ell_{K}>0$ such that
    $|f(x,u)-f(y,u)|\leq\ell_{K}|x-y|$
    for all $x,y\in K$ and each $u\in U$.
Remark 2.14 (Concerning Hypotheses (H)).
Let it be noted that since we restrict our attention to compact control sets, our assumptions encompass standard controlled linear systems of the form
$f(x,u):=Ax+Bu$
where $U\subset\mathbb{R}^{n}$ is compact and $A\in\mathbb{R}^{d\times d}$, $B\in\mathbb{R}^{d\times n}$ are matrices. We also stress that all our results could be extended to unbounded time intervals e.g. by assuming that the dynamics is stable, in the sense that solutions starting from a compact set of initial data remain inside a (possibly larger) compact set.
Throughout our developments, we will often use the shorthand notation $f_{u}\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ for the vector field $x\in\mathbb{R}^{d}\mapsto f(x,u)\in\mathbb{R}^{d}$ associated with some control input $u\in U$, and denote by
$\pazocal{F}:=\Big{\{}f_{u}\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}$
the set of all admissible fields of the control system. In the following proposition, we show that the latter is compact under our working assumptions, a fact that will prove very useful in the sequel.
Proposition 2.15 (Compactness of the set of admissible vector fields).
Under Hypotheses (H), the set $\pazocal{F}\subset C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ is compact for the topology of local uniform convergence.
Proof.
Let $(f_{n})\subset\pazocal{F}$ be a sequence of admissible vector fields. By construction, there exists a corresponding sequence of controls $(u_{n})\subset U$ such that $f_{n}=f_{u_{n}}$ for each $n\geq 1$. Since $(U,\textnormal{{d}}_{U}(\cdot,\cdot))$ is a compact metric space, by assumption, there exists an element $u\in U$ for which
$\textnormal{{d}}_{U}(u_{n},u)\leavevmode\nobreak\ \underset{n\to+\infty}{\longrightarrow}\leavevmode\nobreak\ 0$
along a subsequence $(u_{n})\subset U$ that we do not relabel. Thus, it follows from Hypothesis (H)-$(i)$ that
[…]
for all $x\in\mathbb{R}^{d}$. Besides, it stems from Hypothesis (H)-$(ii)$ that the maps $f_{u_{n}}:K\to\mathbb{R}^{d}$ are uniformly equicontinuous over each compact set $K\subset\mathbb{R}^{d}$, so that the convergence must be uniform, i.e.
[…]
The latter identity being valid for all compact set, this concludes the proof of our claim. ∎
In the following theorem, we recall a standard Cauchy-Lipschitz well-posedness result for the controlled dynamics (2.5), along with elementary estimates satisfied by the underlying flow maps. We point the interested reader to [10, Chapter 2] for detailed proofs thereof.
Theorem 2.16 (Well-posedness, stability and representation of solutions to (2.5)).
Let $(\tau,x)\in[0,T]\times\mathbb{R}^{d}$ be given and suppose that Hypotheses (H) hold. Then for each $u(\cdot)\in\pazocal{U}$, the controlled dynamics (2.5) admits a unique solution $x_{u}(\cdot)\in\textnormal{Lip}([0,T],\mathbb{R}^{d})$ which can be represented explicitly as
$x_{u}(t)=\Phi_{(\tau,t)}^{u}(x)$ (2.6)
for all times $t\in[0,T]$. Therein, the family of maps $(\Phi_{(\tau,t)}^{u})_{\tau,t\in[0,T]}\subset C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ are the flows of homeomorphisms defined as the unique solution of the Cauchy problem
$\Phi_{(\tau,t)}^{u}(x)=x+\int_{\tau}^{t}f\Big{(}\Phi_{(\tau,s)}^{u}(x),u(s)\Big{)}\textnormal{d}s.$ (2.7)
In addition, for each $R>0$, the flows comply with the bound
$\big{|}\Phi_{(\tau,t)}^{u}(x)\big{|}\leq M_{R}$ (2.8)
for all times $\tau,t\in[0,T]$ and every $x\in B(0,R)$, as well as with the uniform Lipschitz estimate
$\big{|}\Phi_{(\tau_{1},t_{1})}^{u}(x)-\Phi_{(\tau_{2},t_{2})}^{u}(y)\big{|}\leq L_{R}\Big{(}|\tau_{1}-\tau_{2}|+|t_{1}-t_{2}|+|x-y|\Big{)},$ (2.9)
for all times $\tau_{1},\tau_{2},t_{1},t_{2}\in[0,T]$ and every $x,y\in B(0,R)$, where $M_{R},L_{R}>0$ are constants which only depend on the magnitudes of $m,T$ and $R$.
We close this section by recollecting a folklore result stating that flow solutions of (2.7) depend continuously on their control input. Therein and in what follows, we will use the observation that $\pazocal{U}$ is a complete metric space when endowed with the distance
$\textnormal{{d}}_{\pazocal{U}}(u(\cdot),v(\cdot)):=\int_{0}^{T}\textnormal{{d}}_{U}(u(t),v(t))\textnormal{d}t$
defined for all $u(\cdot),v(\cdot)\in\pazocal{U}$.
Corollary 2.17 (Continuous dependence in the control).
Under Hypotheses (H), the functional
$u(\cdot)\in\pazocal{U}\mapsto\Phi_{(\cdot,\cdot)}^{u}(\cdot)\in C^{0}([0,T]\times[0,T]\times\mathbb{R}^{d},\mathbb{R}^{d})$
which to a control signal associates the corresponding flow map is continuous.
Proof.
Being somewhat standard and technical, the proof of this result is deferred to Appendix A. ∎
3 Set-valued Koopman operators
In this section, we introduce the set-valued Koopman operators associated with a control system, and discuss some of their fundamental properties. Throughout the section, we will assume that Hypotheses (H) hold and carry out our analysis over the separable normed space $(\pazocal{X},\parallel\cdot\parallel_{\pazocal{X}}):=(C^{0}_{c}(\mathbb{R}^{d},\mathbb{C}),\left\|\cdot\right\|_{C^{0}(\mathbb{R}^{d},\mathbb{C})})$ of continuous and complex-valued observables with compact support.
Definition 3.1 (Set-valued Koopman operators).
We define the set-valued Koopman operators associated with the control system (2.5) by
$\pazocal{K}_{(\tau,t)}:\varphi\in\pazocal{X}\rightrightarrows\Big{\{}\varphi\circ\Phi_{(\tau,t)}^{u}\leavevmode\nobreak\ \,\textnormal{s.t. $u(\cdot)\in\pazocal{U}$}\Big{\}}\subset\pazocal{X}$ (3.1)
for all times $\tau,t\in[0,T]$.
As highlighted in the introduction, the action of the set-valued Koopman operator can be thought of as the reachable set of observables starting from $\varphi\in\pazocal{X}$ and generated by all the admissible control signals. Before investigating these objects further, a few remarks are in order. First, notice that when the dynamics is uncontrolled, i.e. when there is a unique flow $(\Phi_{t})_{t\in[0,T]}\subset C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$, then
$\pazocal{K}_{(\tau,t)}(\varphi)=\{\varphi\circ\Phi_{t-\tau}\}$
and the object introduced in Definition 3.1 coincides with the classical Koopman operator. Besides, it follows from (3.1) together with the fact that flows of nonautonomous dynamical systems form a semigroup for the composition operation that the latter satisfies the identity
$\pazocal{K}_{(\tau,t)}(\varphi)=\pazocal{K}_{(s,t)}\circ\pazocal{K}_{(\tau,s)}(\varphi)$
for all times $\tau,s,t\in[0,T]$, where we implicitly used the natural notation overloading
$\pazocal{K}_{(\tau,t)}(\bm{\Phi}):=\bigcup_{\varphi\in\bm{\Phi}}\pazocal{K}_{(\tau,t)}(\varphi)$
for some set $\bm{\Phi}\subset\pazocal{X}$. Another direct observation that one can make is that $\pazocal{K}_{(\tau,t)}:\pazocal{X}\rightrightarrows\pazocal{X}$ is homogeneous, namely
$\pazocal{K}_{(\tau,t)}(\alpha\varphi)=\alpha\pazocal{K}_{(\tau,t)}(\varphi)$
for each $\alpha\in\mathbb{R}$, as well as subadditive, since
$\pazocal{K}_{(\tau,t)}\big{(}\varphi_{1}+\varphi_{2}\big{)}$ =$=\Big{\{}\big{(}\varphi_{1}+\varphi_{2}\big{)}\circ\Phi_{(\tau,t)}^{u}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u(\cdot)\in\pazocal{U}\Big{\}}$
$\subset\Big{\{}\varphi_{1}\circ\Phi_{(\tau,t)}^{u_{1}}+\varphi_{2}\circ\Phi_{(\tau,t)}^{u_{2}}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u_{1}(\cdot),u_{2}(\cdot)\in\pazocal{U}\Big{\}}=\pazocal{K}_{(\tau,t)}(\varphi_{1})+\pazocal{K}_{(\tau,t)}(\varphi_{2})$
for each $\varphi_{1},\varphi_{2}\in\pazocal{X}$. However, the sets $\pazocal{K}_{(\tau,t)}(\varphi)$ are a priori not convex in general, and the Koopman operators are thus not linear in the set-valued sense discussed in Definition 2.12. We shall see nonetheless that they still enjoy nice characterisations for their adjoints along with a meaningful variant of the spectral mapping theorem relating the eigenvalues of the semigroup and its generator.
Remark 3.2 (Set-valued Koopman operators for differential inclusions).
We would like to stress that while we define set-valued Koopman operators for controlled systems, most of our results remain valid for dynamical systems modelled more generally by differential inclusions, that is
$\dot{x}(t)\in F(x(t))$ (3.2)
with $F:\mathbb{R}^{d}\rightrightarrows\mathbb{R}^{d}$ being a Lipschitz continuous set-valued mapping with nonempty compact images. In this case, one can simply define the set-valued Koopman operators as
[…]
for all times $t\in[0,T]$. In the situation where $F:\mathbb{R}^{d}\rightrightarrows\mathbb{R}^{d}$ has convex images – an assumption that is quite standard in control theory, see e.g. [5, 15, 50], and covered by Hypothesis (C) below –, the Lipschitz parametrisation theorem of [5, Theorem 9.7.1] yields the existence of a compact set $U\subset\mathbb{R}^{d}$ along with a Lipschitz map $f:\mathbb{R}^{d}\times U\to\mathbb{R}^{d}$ such that $F(x)=\{f(x,u)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\}$ for all $x\in\mathbb{R}^{d}$.
Remark 3.3 (Concerning time-varying systems).
We also mention that the majority of our developments would still hold, up to some technical adjustments, for time-varying vector fields $(t,x,u)\in[0,T]\times\mathbb{R}^{d}\times U\mapsto f(t,x,u)\in\mathbb{R}^{d}$ which are $\mathscr{L}^{1}$-measurable in $t\in[0,T]$. That being said, we chose to stick to the simpler framework of time-invariant control systems for the sake of readability, and point the interested reader to [21] for some insight on how our results could be rephrased in this context.
In the following proposition, we establish basic regularity results for the set-valued Koopman operators, akin to those already known in the context of classical dynamical systems.
Proposition 3.4 (Regularity properties of set-valued Koopman operators).
Under Hypotheses (H), the set-valued Koopman operator $\varphi\in\pazocal{X}\rightrightarrows\pazocal{K}_{(\tau,t)}(\varphi)$ is $1$-Lipschitz for all times $\tau,t\in[0,T]$, and the set-valued map $(\tau,t)\in[0,T]\times[0,T]\rightrightarrows\pazocal{K}_{(\tau,t)}(\varphi)$ is continuous for each $\varphi\in\pazocal{X}$.
Proof.
We start by studying the regularity of $\pazocal{K}_{(\tau,t)}:\pazocal{X}\rightrightarrows\pazocal{X}$ for some given $\tau,t\in[0,T]$. We fix an arbitrary pair $\varphi_{1},\varphi_{2}\in\pazocal{X}$, and note that for every element $\psi_{(\tau,t)}^{1}\in\pazocal{K}_{(\tau,t)}(\varphi_{1})$, there exists some $u_{1}(\cdot)\in\pazocal{U}$ such that $\psi_{(\tau,t)}^{1}=\varphi_{1}\circ\Phi_{(\tau,t)}^{u_{1}}$. Then, recalling that $(\pazocal{X},\parallel\cdot\parallel_{\pazocal{X}})$ is a vector space, one has that
$\psi_{(\tau,t)}^{1}$ $=\varphi_{1}\circ\Phi_{(\tau,t)}^{u_{1}}$
$=\varphi_{2}\circ\Phi_{(\tau,t)}^{u_{1}}+\big{(}\varphi_{2}-\varphi_{1}\big{)}\circ\Phi_{(\tau,t)}^{u_{1}}$
$\in\pazocal{K}_{(\tau,t)}(\varphi_{2})+\mathbb{B}_{\pazocal{X}}\big{(}0,\parallel\varphi_{1}-\varphi_{2}\parallel_{\pazocal{X}}\big{)},$
which implies that $\pazocal{K}_{(\tau,t)}(\varphi_{1})\subset\pazocal{K}_{(\tau,t)}(\varphi_{2})+\mathbb{B}_{\pazocal{X}}\big{(}0,\parallel\varphi_{1}-\varphi_{2}\parallel_{\pazocal{X}})$ and is therefore tantamount to the $1$-Lipschitzianity of $\pazocal{K}_{(\tau,t)}:\pazocal{X}\rightrightarrows\pazocal{X}$ for all times $\tau,t\in[0,T]$. Consider now an element $\varphi\in\pazocal{X}$, and observe that owing to the uniform bound displayed in (2.8) of Theorem 2.16, the set
$K_{\varphi}:=\overline{\bigcup_{u(\cdot)\in\pazocal{U}}\bigcup_{\tau,t\in[0,T]}\Phi_{(t,\tau)}^{u}(\textnormal{supp}(\varphi))}\subset\mathbb{R}^{d}$
is well-defined and compact. Let $\tau_{1},t_{1}\in[0,T]$ be arbitrary, fix some $\psi_{(\tau_{1},t_{1})}\in\pazocal{K}_{(\tau_{1},t_{1})}(\varphi)$, and note that there exists a control $u_{1}(\cdot)\in\pazocal{U}$ such that $\psi_{(\tau_{1},t_{1})}=\varphi\circ\Phi_{(\tau_{1},t_{1})}^{u_{1}}$. It then follows from the flow regularity estimate (2.9) that for any $\varepsilon>0$, there exists some $\delta>0$ such that
$\sup_{x\in K_{\varphi}}|\Phi_{(\tau_{1},t_{1})}^{u_{1}}(x)-\Phi^{u_{1}}_{(\tau,t)}(x)|\leq\varepsilon$
whenever $\tau,t\in[0,T]$ satisfy $(|\tau_{1}-\tau|+|t_{1}-t|)\leq\delta$. Observing in turn that $\varphi\in\pazocal{X}$ is uniformly continuous over its support while letting $\psi_{(\tau,t)}:=\varphi\circ\Phi_{(\tau,t)}^{u_{1}}\in\pazocal{K}_{(\tau,t)}(\varphi)$, it then follows up to potentially choosing a smaller $\delta>0$ that
$\parallel\psi_{(\tau_{1},t_{1})}-\psi_{(\tau,t)}\parallel_{\pazocal{X}}\;=\sup_{x\in K_{\varphi}}\big{|}\varphi\circ\Phi_{(\tau_{1},t_{1})}^{u_{1}}(x)-\varphi\circ\Phi_{(\tau,t)}^{u_{1}}(x)\big{|}\leavevmode\nobreak\ \leq\,\varepsilon.$
Thence, the set-valued map $(\tau,t)\in[0,T]\times[0,T]\rightrightarrows\pazocal{K}_{(\tau,t)}(\varphi)\subset\pazocal{X}$ is lower-semicontinuous at $(\tau_{1},t_{1})\in[0,T]\times[0,T]$ for every such pair, and thus over the whole time interval. It can be shown via very similar arguments that it is also upper-semicontinuous, which ends the proof of our claim. ∎
In what follows, we build on the previous result to prove a simple, yet enlightening representation formula for time-dependent Koopman observables.
Proposition 3.5 (Representation formula for curves of Koopman observables).
Given some observalbe $\varphi\in\pazocal{X}$, a function $(\tau,t)\in[0,T]\times[0,T]\to\psi_{(\tau,t)}\in\pazocal{X}$ is a measurable selection in the set-valued map $(\tau,t)\in[0,T]\times[0,T]\rightrightarrows\pazocal{K}_{(\tau,t)}(\varphi)$ if and only if there exists an $\mathscr{L}^{1}\times\mathscr{L}^{1}$-measurable map $(\tau,t)\in[0,T]\times[0,T]\mapsto u_{\tau,t}(\cdot)\in\pazocal{U}$ such that
$\psi_{(\tau,t)}=\varphi\circ\Phi_{(\tau,t)}^{u_{\tau,t}}$
for $\mathscr{L}^{1}\times\mathscr{L}^{1}$-almost every $(\tau,t)\in[0,T]\times[0,T]$.
Proof of Proposition 3.5.
Clearly, if $\psi_{(\tau,t)}=\varphi\circ\Phi_{(\tau,t)}^{u_{\tau,t}}$ for some measurable map $(\tau,t)\in[0,T]\times[0,T]\mapsto u_{\tau,t}(\cdot)\in\pazocal{U}$, then $\psi_{(\tau,t)}\in\pazocal{K}_{(\tau,t)}(\varphi)$ for $\mathscr{L}^{1}\times\mathscr{L}^{1}$-almost every $(\tau,t)\in[0,T]\times[0,T]$. Moreover, the map $(\tau,t)\in[0,T]\times[0,T]\mapsto\psi_{(\tau,t)}\in\pazocal{X}$ is then measurable by the continuity of
$(\tau,t,u(\cdot))\in[0,T]\times[0,T]\times\pazocal{U}\mapsto\Phi_{(\tau,t)}^{u}\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$
that stems from Theorem 2.16 and Corollary 2.17. Suppose now that $(\tau,t)\in[0,T]\times[0,T]\mapsto\psi_{(\tau,t)}\in\pazocal{K}_{(\tau,t)}(\varphi)$ is an arbitrary measurable selection in the set-valued Koopman operator, and notice that the latter can be expressed as
$\pazocal{K}_{(\tau,t)}(\varphi)=\Big{\{}\Psi(\tau,t,u(\cdot))\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u(\cdot)\in\pazocal{U}\Big{\}}$
where $\Psi:(\tau,t,u(\cdot))\in[0,T]\times[0,T]\times\pazocal{U}\mapsto\varphi\circ\Phi_{(\tau,t)}^{u}\in\pazocal{X}$ is continuous, again as a consequence of Theorem 2.16 and Corollary 2.17. Because the completion of $(\pazocal{X},\parallel\cdot\parallel_{\pazocal{X}})$ is a separable Banach space, we may apply Filippov’s selection theorem to recover the existence of a measurable map $(\tau,t)\in[0,T]\times[0,T]\mapsto u_{\tau,t}(\cdot)\in\pazocal{U}$ such that
$\psi_{(\tau,t)}=\Psi(\tau,t,u_{\tau,t}(\cdot))=\varphi\circ\Phi_{(\tau,t)}^{u_{\tau,t}}$
for $\mathscr{L}^{1}\times\mathscr{L}^{1}$-almost every $\tau,t\in[0,T]$, which closes the proof. ∎
Remark 3.6 (On the representation of Koopman observables).
In essence, the previous result shows that while every element in $\pazocal{K}_{(\tau,t)}(\varphi)$ can be expressed using of a single controlled flow when both $\tau,t\in[0,T]$ are fixed, a time-dependent family $(\tau,t)\in[0,T]\times[0,T]\mapsto\psi_{(\tau,t)}\in\pazocal{K}_{(\tau,t)}(\varphi)$ may jump between the realisations of different controlled dynamics, as illustrated by Figure 3.1.
[…]
Figure 3.1: Representation of a particular curve of Koopman observables corresponding to measurements evaluated along a single controlled flow (left) and of a general curve of observables that may a priori jump in a measurable way between different controlled flows (right).
Throughout the manuscript, many of our results will rely crucially on the assumption that the set of controlled velocities is convex in a way which faithfully captures the fact that Koopman operators act globally – instead of pointwisely – on the underlying state space.
Hypothesis (C).
The set of controlled vector fields $\pazocal{F}:=\big{\{}f_{u}\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\big{\}}$ is convex.
A prototypical example of controlled velocity set satisfying Hypotheses (H) and (C) above is given by nonlinear control-affine systems of the form
$f_{u}(x):=f_{0}(x)+\sum_{k=1}^{n}u_{k}f_{k}(x),$
where $U\subset\mathbb{R}^{n}$ is convex and $f_{0},\dots,f_{n}\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ are locally Lipschitz vector fields with sublinear growth. It should be stressed however that Hypothesis (C) is in general much stronger than the usual condition that $\{f(x,u)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\}\subset\mathbb{R}^{d}$ be a convex set for all fixed $x\in\mathbb{R}^{d}$.
In the ensuing proposition, we establish the compactness of the images of the Koopman operators, which stems from the functional convexity of the set of admissible vector fields.
Proposition 3.7 (Topological properties of Koopman operators).
Suppose that Hypotheses (H) and (C) hold. Then, the sets $\pazocal{K}_{(\tau,t)}(\varphi)\subset\pazocal{X}$ are compact for all times $\tau,t\in[0,T]$ and each $\varphi\in\pazocal{X}$, and the operators $\pazocal{K}_{(\tau,t)}:\pazocal{X}\rightrightarrows\pazocal{X}$ are closed processes.
Proof.
To show that, given $\tau,t\in[0,T]$ and some $\varphi\in\pazocal{X}$, the sets $\pazocal{K}_{(\tau,t)}(\varphi)$ are compact, fix $(\psi_{(\tau,t)}^{n})\subset\pazocal{K}_{(\tau,t)}(\varphi)$ and let $(u_{n}(\cdot))\subset\pazocal{U}$ be the sequence of signals such that $\psi_{(\tau,t)}^{n}=\varphi\circ\Phi_{(\tau,t)}^{u_{n}}$ for each $n\geq 1$. Recall also that by Theorem 2.16, there exist for each $R>0$ a pair of constants $M_{R},L_{R}>0$ such that
$\big{|}\Phi_{(\tau,t)}^{u_{n}}(x)\big{|}\leq M_{R}\qquad\text{and}\qquad\big{|}\Phi_{(\tau,t)}^{u_{n}}(x)-\Phi_{(\tau,t)}^{u_{n}}(y)\big{|}\leq L_{R}|x-y|,$ (3.3)
for $x,y\in B(0,R)$ and each $n\geq 1$. Therefore, the sequence of maps $(\Phi_{(\tau,t)}^{u_{n}})\subset C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ is locally uniformly valued in a compact set and locally equi-Lipschitz, which by the Ascoli-Arzelà theorem (see e.g. [27, Chapter 7 Theorem 18]) yields the existence of an element $\Phi_{(\tau,t)}\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ such that
[…]
for each compact set $K\subset\mathbb{R}^{d}$, along a subsequence $(u_{n_{k}}(\cdot))\subset\pazocal{U}$. Observing now that the set
$K_{\varphi}:=\overline{\bigcup_{n\geq 1}\bigcup_{\tau,t\in[0,T]}\Phi^{u_{n}}_{(t,\tau)}(\textnormal{supp}(\varphi))}$
is well-defined and compact as a consequence of (3.3), it necessarily follows that
[…]
along the same subsequence. Hence, to close the argument, there remains to show that
$\Phi_{(\tau,t)}=\Phi_{(\tau,t)}^{u}$ (3.5)
for some admissible control $u(\cdot)\in\pazocal{U}$. To this end, recall first that by Proposition 2.15, the set of admissible velocities $\pazocal{F}\subset C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ is convex and compact for the topology of local uniform convergence. Thus, for each compact set $K\subset\mathbb{R}^{d}$, the collection of restricted vector fields
$\pazocal{F}_{K}:=\Big{\{}f_{|K}\in C^{0}(K,\mathbb{R}^{d})\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ f\in\pazocal{F}\Big{\}}$
is a compact and convex set as well. Therefore, we can inductively apply the compactness criterion of Proposition 2.3 on a countably increasing family of balls covering $\mathbb{R}^{d}$ along with a diagonal argument to obtain the existence of an $\mathscr{L}^{1}$-measurable map $s\in[0,T]\mapsto f(s)\in\pazocal{F}$ satisfying
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\zeta\cdot\Big{(}f(s,x)-f_{u_{n_{k}}(s)}(x)\Big{)}\textnormal{d}\mu(s)(x)\textnormal{d}s\leavevmode\nobreak\ \underset{k\to+\infty}{\longrightarrow}\leavevmode\nobreak\ 0$ (3.6)
for every $\mu:[0,T]\to\mathscr{M}(K,\mathbb{C})$ complying with the conditions of Proposition 2.3 for $\pazocal{X}:=C^{0}(K,\mathbb{R}^{d})$, each $\zeta\in\mathbb{R}^{d}$ and every compact set $K\subset\mathbb{R}^{d}$. Therein, the convergence holds along a subsequence that may depend on the set $K\subset\mathbb{R}^{d}$ itself, and that we do not relabel. At this stage, we consider the set
$K_{x}:=\overline{\bigcup_{n\geq 1}\bigcup_{s\in[0,T]}\Phi^{u_{n}}_{(\tau,s)}(x)},$
defined for any $x\in\mathbb{R}^{d}$, and note that it is compact by (3.3). Choosing now the particular measure-valued map
$\mu(s):=\mathds{1}_{[\tau,t]}(s)\,\delta_{\Phi_{(\tau,s)}(x)}\in\mathscr{M}(K_{x},\mathbb{R})$
defined for $\mathscr{L}^{1}$-almost every $s\in[0,T]$, it follows from (3.6) that
$\int_{\tau}^{t}\zeta\cdot\Big{(}f(s,\Phi_{(\tau,s)}(x))-f_{u_{n_{k}}(s)}\big{(}\Phi_{(\tau,s)}(x)\big{)}\Big{)}\textnormal{d}s\leavevmode\nobreak\ \underset{k\to+\infty}{\longrightarrow}\leavevmode\nobreak\ 0$
for every $\zeta\in\mathbb{R}^{d}$, along a subsequence that may depend on $(t,x)\in[0,T]\times\mathbb{R}^{d}$. Besides, observing that
$\int_{\tau}^{t}\,\left\|f_{u_{n_{k}}(s)}\big{(}\Phi_{(\tau,s)}\big{)}-f_{u_{n_{k}}(s)}\big{(}\Phi_{(\tau,s)}^{u_{n_{k}}}\big{)}\right\|_{C^{0}(K_{x},\mathbb{R}^{d})}\textnormal{d}s\leavevmode\nobreak\ \underset{k\to+\infty}{\longrightarrow}\leavevmode\nobreak\ 0$
as a byproduct of (3.4) and Hypothesis (H)-$(ii)$, we can conclude by merging the previous two equations that
$\int_{0}^{t}f_{u_{n_{k}}(s)}\Big{(}\Phi_{(\tau,s)}^{u_{n_{k}}}(x)\Big{)}\textnormal{d}s\leavevmode\nobreak\ \underset{k\to+\infty}{\longrightarrow}\leavevmode\nobreak\ \int_{0}^{t}f(s,\Phi_{(\tau,s)}(x))\textnormal{d}s$ (3.7)
along a subsequence that may depend on $(t,x)\in[0,T]\times\mathbb{R}^{d}$. Since the latter are arbitrary, we can deduce from (3.4) and (3.7) that
$\Phi_{(\tau,t)}(x)=x+\int_{\tau}^{t}f(s,\Phi_{(\tau,s)}(x))\textnormal{d}s$ (3.8)
for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$. Finally, noting that
$f(s)\in\pazocal{F}=\Big{\{}f_{u}\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}$
for $\mathscr{L}^{1}$-almost every $s\in[0,T]$, while observing that the map $u\in U\mapsto f_{u}\in C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ is continuous under Hypotheses (H), one may apply Theorem 2.10 to infer the existence of some $u(\cdot)\in\pazocal{U}$ such that
$f(s)=f_{u(s)}$
for $\mathscr{L}^{1}$-almost every $s\in[0,T]$. This, together with (3.8), is tantamount to (3.5) and thus concludes the proof of our first claim.
We now prove that $\pazocal{K}_{(\tau,t)}:\pazocal{X}\rightrightarrows\pazocal{X}$ is a closed process for all times $\tau,t\in[0,T]$, namely that its graph is a closed cone. The latter of these properties is straightforward, since for each $\alpha>0$ and every $(\varphi,\psi_{(\tau,t)})\in\textnormal{Graph}(\pazocal{K}_{(\tau,t)})$, one has that
$\alpha\psi_{(\tau,,t)}=\alpha\varphi\circ\Phi_{(\tau,t)}^{u}\in\pazocal{K}_{(\tau,t)}(\alpha\varphi)$
where $u(\cdot)\in\pazocal{U}$. In order to prove that $\textnormal{Graph}(\pazocal{K}_{(\tau,t)})$ is closed, we consider a sequence $((\varphi_{n},\psi_{(\tau,t)}^{n}))\subset\textnormal{Graph}(\pazocal{K}_{(\tau,t)})$ such that
$\parallel\varphi-\varphi_{n}\parallel_{\pazocal{X}}\underset{n\to+\infty}{\longrightarrow}0\qquad\text{and}\qquad\parallel\psi-\psi_{(\tau,t)}^{n}\parallel_{\pazocal{X}}\underset{n\to+\infty}{\longrightarrow}0,$ (3.9)
and recall that by definition, there exists $(u_{n}(\cdot))\subset\pazocal{U}$ so that
$\psi_{(\tau,t)}^{n}=\varphi_{n}\circ\Phi_{(\tau,t)}^{u_{n}}$
for each $n\geq 1$. Then, upon repeating the compactness argument detailed hereinabove combined with (3.9), it follows that $\psi\in\pazocal{K}_{(\tau,t)}(\varphi)$, which concludes the proof. ∎
4 Set-valued Liouville and Perron-Frobenius operators
In this section, we define the set-valued counterparts of the Liouville and Perron-Frobenius operators, for which we provide several structure results and explicit characterisations. Throughout what follows, we denote by $(\pazocal{D},\parallel\cdot\parallel_{\pazocal{D}}):=(C^{1}_{c}(\mathbb{R}^{d},\mathbb{C}),\parallel\cdot\parallel_{C^{1}(\mathbb{R}^{d},\mathbb{C})})$ the separable normed space of continuously differentiable complexed-valued functions with compact support.
4.1 Set-valued Liouville operators and Koopman dynamics
In what ensues, we define the set-valued Liouville operator as the collection of all Liouville operators associated with each individual controls, and show that the latter can be rigorously related to the infinitesimal behaviour of the set-valued Koopman semigroup $(\pazocal{K}_{(\tau,t)})_{\tau,t\in[0,T]}$ at time $t=\tau$. These results provide theoretical underpinning to the works defining the Koopman and Liouville operators for control systems as the collections of operators associated to each control input, see e.g. [39, 40].
Definition 4.1 (Set-valued Liouville operators).
We define the set-valued Liouville operator $\pazocal{L}:\pazocal{D}\rightrightarrows\pazocal{X}$ associated with (2.5) as
$\pazocal{L}(\varphi):=\Big{\{}\nabla_{x}\varphi\cdot f_{u}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}\subset\pazocal{X}$ (4.1)
for each $\varphi\in\pazocal{D}$.
We start by establishing some simple topological properties of these Liouville operators. These results will prove insightful when studying its point spectrum further down in Section 4.3.
Proposition 4.2 (Topological properties of the Liouville operator).
Suppose that Hypotheses (H) and (C) hold. Then, the set-valued Liouville operator $\pazocal{L}:\pazocal{D}\rightrightarrows\pazocal{X}$ is a fan with compact images.
Proof.
The fact that $\pazocal{L}:\pazocal{D}\rightrightarrows\pazocal{X}$ has convex and compact images is a direct consequence of Hypothesis (C), on the one hand, and of Hypotheses (H) together with Proposition 2.15 on the other hand. Besides, it follows from the very definition of the set-valued Liouville operator that $0\in\pazocal{L}(0)$ and
$\pazocal{L}(\varphi_{1}+\varphi_{2})$ =$=\Big{\{}\big{(}\nabla_{x}\varphi_{1}+\nabla_{x}\varphi_{2}\big{)}\cdot f_{u}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}$
$\subset\Big{\{}\nabla_{x}\varphi_{1}\cdot f_{u_{1}}+\nabla_{x}\varphi_{2}\cdot f_{u_{2}}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u_{1},u_{2}\in U\Big{\}}=\pazocal{L}(\varphi_{1})+\pazocal{L}(\varphi_{2})$
for each $\varphi_{1},\varphi_{2}\in\pazocal{D}$. Hence, there simply remains to show that $\pazocal{L}:\pazocal{D}\rightrightarrows\pazocal{X}$ is a closed process, namely that its graph is a closed cone. The latter of these properties is straightforward, as for each $\alpha>0$ and every $(\varphi,\psi)\in\textnormal{Graph}(\pazocal{L})$, it clearly holds that
$\alpha\psi=\alpha\nabla_{x}\varphi\cdot f_{u}\in\pazocal{L}(\alpha\varphi)$
where $u\in U$ is some fixed control value. Suppose now that we are given a sequence $((\varphi_{n},\psi_{n}))\subset\textnormal{Graph}(\pazocal{L})$ such that
$\parallel\varphi-\varphi_{n}\parallel_{\pazocal{D}}\leavevmode\nobreak\ \underset{n\to+\infty}{\longrightarrow}\leavevmode\nobreak\ 0\qquad\text{and}\qquad\parallel\psi-\psi_{n}\parallel_{\pazocal{X}}\leavevmode\nobreak\ \underset{n\to+\infty}{\longrightarrow}\leavevmode\nobreak\ 0,$ (4.2)
for some $(\varphi,\psi)\in\pazocal{D}\times\pazocal{X}$. This fact together with the definition (4.1) of the Liouville operator yield the existence of a sequence $(u_{n})\subset U$ for which
$\psi_{n}=\nabla_{x}\varphi_{n}\cdot f_{u_{n}}$
for all $n\geq 1$. The conclusion simply follows then from the compactness of the set of admissible velocities established in Proposition 2.15. ∎
In the following theorem, we prove a result which, in the spirit of [22, Section 2], establishes a connection between the Liouville operator and set-valued derivatives of the Koopman semigroup.
Theorem 4.3 (Set-valued Liouville operators as generators of the Koopman semigroups).
Suppose that Hypotheses (H) hold and fix a time $\tau\in[0,T]$. Then, the following inclusions
$\pazocal{L}(\varphi)\subset\underset{\leavevmode\nobreak\ t\to\tau}{\textnormal{Liminf}}\,\frac{\pazocal{K}_{(\tau,t)}(\varphi)-\varphi}{t-\tau}$ (4.3)
and
[…]
are satisfied for each $\varphi\in\pazocal{D}$. Furthermore, if Hypothesis (C) holds, then the set-valued Liouville operator $\pazocal{L}:\pazocal{D}\rightrightarrows\pazocal{X}$ is the infinitesimal generator of the Koopman semigroup, in the sense that
$\pazocal{L}(\varphi)=\underset{\leavevmode\nobreak\ t\to\tau}{\textnormal{Lim}}\frac{\pazocal{K}_{(\tau,t)}(\varphi)-\varphi}{t-\tau}$ (4.5)
for each $\varphi\in\pazocal{D}$.
Proof.
It is quite clear from the definition of lower and upper limits of sequences of sets provided in Section 2.2 above that
$\underset{\leavevmode\nobreak\ t\to\tau}{\textnormal{Liminf}}\,\frac{\pazocal{K}_{(\tau,t)}(\varphi)-\varphi}{t-\tau}\subset\underset{\leavevmode\nobreak\ t\to\tau}{\textnormal{Limsup}}\,\frac{\pazocal{K}_{(\tau,t)}(\varphi)-\varphi}{t-\tau}$
for each $\varphi\in\pazocal{D}$. Remarking thus that under Hypothesis (C), there holds
$\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{L}(\varphi)=\overline{\textnormal{co}}\hskip 1.42271pt\Big{\{}\nabla_{x}\varphi\cdot f\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ f\in\pazocal{F}\Big{\}}=\Big{\{}\nabla_{x}\varphi\cdot f\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ f\in\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{F}\Big{\}}=\pazocal{L}(\varphi),$
it would directly follow from (4.3) and (4.4) that the set-valued Liouville operator coincides with the derivative displayed in (4.5).
We thus start by establishing the liminf inclusion (4.3), which amounts to finding for each element $\bar{u}\in U$ a curve of Koopman observables $t\in[0,T]\mapsto\psi_{(\tau,t)}\in\pazocal{K}_{(\tau,t)}(\varphi)$ such that
$\lim_{t\to\tau}\frac{\psi_{(\tau,t)}-\varphi}{t-\tau}=\nabla_{x}\varphi\cdot f_{\bar{u}}.$ (4.6)
To this end, consider the constant control signal given by $u(t):=\bar{u}$ for all times $t\in[0,T]$, and let $\psi_{(\tau,t)}:=\varphi\in\Phi_{(\tau,t)}^{u}$. Observe then that, by construction, one has that
$\Phi_{(\tau,t)}^{u}(x)$ $=x+\int_{\tau}^{t}f\Big{(}\Phi_{(\tau,s)}^{u}(x),\bar{u}\Big{)}\textnormal{d}s$
$=x+(t-\tau)f_{\bar{u}}(x)+\int_{\tau}^{t}\bigg{(}f_{\bar{u}}\big{(}\Phi_{(\tau,s)}^{u}(x)\big{)}-f_{\bar{u}}(x)\bigg{)}\textnormal{d}s,$
for all $x\in\mathbb{R}^{d}$. By combining Hypothesis (H)-$(ii)$ and the estimate of (2.9) in Theorem 2.16 within the latter identity, one may further obtain
$\sup_{x\in K_{\varphi}}\big{|}\Phi_{(\tau,t)}^{u}(x)-x-(t-\tau)f_{\bar{u}}(x)\big{|}\leq C_{\varphi}|t-\tau|^{2}$ (4.7)
for all times $t\in[0,T]$, where $C_{\varphi}>0$ is a constant and $K_{\varphi}\subset\mathbb{R}^{d}$ is the compact set defined by
$K_{\varphi}:=\overline{\bigcup_{u(\cdot)\in\pazocal{U}}\bigcup_{t\in[0,T]}\Phi_{(t,\tau)}^{u}\big{(}\textnormal{supp}(\varphi)\big{)}},$ (4.8)
whose existence is guaranteed by (2.8). Thence, recalling that $\varphi\in C^{1}_{c}(\mathbb{R}^{d},\mathbb{C})$, it follows from (4.7) combined with (4.8) that
$\lim_{t\to\tau}\Bigg{\|}\frac{\varphi\circ\Phi_{(\tau,t)}^{u}-\varphi-(t-\tau)\nabla_{x}\varphi\cdot f_{\bar{u}}}{t-\tau}\Bigg{\|}_{\pazocal{X}}=0,$
which then yields (4.6) by what precedes, since $\bar{u}\in U$ is arbitrary.
To complete the proof, there remains to derive the limsup inclusion (4.4), which can be equivalently recast as the requirement that
$\lim_{t_{n}\to\tau}\frac{\psi_{(\tau,t_{n})}-\varphi}{t_{n}-\tau}\in\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{L}(\varphi)$ (4.9)
for every pair of sequences $(t_{n})\subset[0,T]\setminus\{\tau\}$ and $(\psi_{(\tau,t_{n})})\subset\pazocal{X}$ along which the aforewritten limit exists and $\psi_{(\tau,t_{n})}\in\pazocal{K}_{(\tau,t_{n})}(\varphi)$ for each $n\geq 1$. Fixing such a pair, it follows from Proposition 3.5 that
$\psi_{(\tau,t_{n})}=\varphi\circ\Phi_{(\tau,t_{n})}^{u_{n}}$
for each $n\geq 1$ and some sequence $(u_{n}(\cdot))\subset\pazocal{U}$. Then, by definition of the characteristic flows
$\Phi_{(\tau,t_{n})}^{u_{n}}(x)$ $=x+\int_{\tau}^{t_{n}}f\Big{(}\Phi_{(\tau,s)}^{u_{n}}(x),u_{n}(s)\Big{)}\textnormal{d}s$ (4.10)
$=x+(t_{n}-\tau)\bigg{(}\frac{1}{t_{n}-\tau}\int_{\tau}^{t_{n}}f(x,u_{n}(s))\textnormal{d}s\bigg{)}+\int_{\tau}^{t_{n}}\bigg{(}f\Big{(}\Phi_{(\tau,s)}^{u_{n}}(x),u_{n}(s)\Big{)}-f(x,u_{n}(s))\bigg{)}\textnormal{d}s$
for each $n\geq 1$. Denoting by $K_{\varphi}\subset\mathbb{R}^{d}$ the compact set associated with $\varphi\in\pazocal{X}$ via (4.8), it holds that
$\sup_{x\in K_{\varphi}}\bigg{|}\int_{\tau}^{t_{n}}\bigg{(}f\Big{(}\Phi_{(\tau,s)}^{u_{n}}(x),u_{n}(s)\Big{)}-f(x,u_{n}(s))\bigg{)}\textnormal{d}s\bigg{|}\leq C_{\varphi}|t_{n}-\tau|^{2},$ (4.11)
for some constant $C_{\varphi}>0$ and each $n\geq 1$. Besides, regarding for any given $n\geq 1$ and each compact set $K\subset\mathbb{R}^{d}$ the map $t\in[0,T]\mapsto f_{u_{n}(t)}\in C^{0}(K,\mathbb{R}^{d})$ as an elements of the Bochner space $L^{1}([0,T],C^{0}(K,\mathbb{R}^{d}))$, it follows from Proposition 2.2 that
$\frac{1}{t_{n}-\tau}\int_{\tau}^{t_{n}}f_{u_{n}(s)}\textnormal{d}s\in\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{F}$
for each $n\geq 1$. Under Hypotheses (H), it can be checked that the mappings
$F_{n}:x\in K_{\varphi}\mapsto\frac{1}{t_{n}-\tau}\int_{\tau}^{t_{n}}f_{u_{n}(s)}(x)\textnormal{d}s\in\mathbb{R}^{d}$
form a uniformly bounded and equi-Lipschitz family in $C^{0}(K_{\varphi},\mathbb{R}^{d})$, which by the standard Ascoli-Arzelà theorem (see e.g. [44, Theorem 11.28]) combined with the fact that $\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{F}\subset C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$ is closed as a consequence of Proposition 2.15, implies that there exists an element $F\in\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{F}$ such that
$\lim_{k\to+\infty}\left\|F-F_{n_{k}}\right\|_{C^{0}(K_{\varphi},\mathbb{R}^{d})}=0$ (4.12)
along a vanishing subsequence $t_{n_{k}}\to\tau$. Owing to the fact that $\varphi\in C^{1}_{c}(\mathbb{R}^{d},\mathbb{C})$, it then follows from (4.10), (4.11) and (4.12) that
$\psi_{(\tau,t_{n_{k}})}=\varphi\circ\Phi^{u_{n_{k}}}_{(\tau,t_{n_{k}})}=\varphi+(t_{n_{k}}-\tau)\nabla_{x}\varphi\cdot F+o(|t_{n_{k}}-\tau|)$
as $t_{n_{k}}\to\tau$. Since we assumed that the limit in (4.9) exists, this last identity ultimately implies that
$\lim_{k\to+\infty}\frac{\psi_{(\tau,t_{n_{k}})}-\varphi}{t_{n_{k}}-\tau}=\nabla_{x}\varphi\cdot F\in\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{L}(\varphi)$
which closes the proof of Theorem 4.3. ∎
Remark 4.4 (On the relation between the set-valued Koopman and Liouville operators).
The previous theorem provides us with two insights on the structure of the Koopman and Liouville operators for control systems. Firstly, the inclusions (4.3) and (4.4) show that, even when the dynamics is not convex-valued, the infinitesimal behaviour of the set-valued Koopman operators is captured by the natural set-valued counterpart of the Liouville operators introduced in Definition 4.1. In Theorem 4.16 below, we shall see how this particular fact allows for the derivation of a set-valued spectral mapping theorem, that can in particular be applied to finite control sets which are inherently non-convex.
Secondly, the identity (4.5) means in turn that, when the admissible velocities are convex in the sense of Hypothesis (C), the set-valued Liouville operator is the infinitesimal generator of the Koopman operator. This transcribes the fact that even though time-dependent Koopman observables cannot be represented in general by a single controlled flow – see Figure 3.1 and Proposition 3.5 above –, the set of all possible infinitesimal evolutions at a time $t$ close to $\tau$ coincides exactly with those generated by a controlled vector field $f_{u}\in\pazocal{F}$ with $u\in U$.
Remark 4.5 (On set-valued derivatives and Liouville operators).
The set-theoretic limit of difference quotient appearing in (4.5) is usually called the adjacent or intermediate derivative in the literature, see e.g. [5, Section 5.2]. This denomination comes from the fact that its graph is the adjacent cone (see e.g. [5, Section 4.1.3]) to that of the set-valued mapping $t\in[0,T]\rightrightarrows\pazocal{K}_{(\tau,t)}(\varphi)\in\pazocal{X}$ at $t=\tau$. While one could alternatively wish to define or characterise $\pazocal{L}:\pazocal{D}\rightrightarrows\pazocal{X}$ by means of other set-valued derivatives, involving e.g. contingent or Clarke cones [5, Section 5.1 and 5.2], adjacent derivatives are known to appear more naturally when investigating the infinitesimal behaviour of reachable sets [22].
We close this section by showing in the next proposition how the set-valued Liouville operator encodes the dynamics of those time-dependent Koopman observables which are generated by precisely one control signal, and thus follow an admissible trajectory of the system.
Proposition 4.6 (Dynamics of time-dependent Koopman observables).
Fix some $\varphi\in\pazocal{D}$, suppose that Hypotheses (H) hold and assume that $f_{u}\in C^{1}(\mathbb{R}^{d},\mathbb{R}^{d})$ for each $u\in U$. Then, a family of Koopman observables $(\tau,t)\in[0,T]\times[0,T]\mapsto\psi_{(\tau,t)}\in\pazocal{K}_{(\tau,t)}(\varphi)$ is of the form $\psi_{(\tau,t)}=\varphi\circ\Phi_{(\tau,t)}^{u}$ for some fixed $u(\cdot)\in\pazocal{U}$ if and only if it is a strong solution of the differential inclusion
[…]
in the space of observables $(\pazocal{X},\parallel\cdot\parallel_{\pazocal{X}})$.
Proof.
We start by assuming that $\psi_{(\tau,t)}=\varphi\circ\Phi_{(\tau,t)}^{u}\in\pazocal{D}$ for some $u(\cdot)\in\pazocal{U}$, and note that by classical results from nonautonomous dynamical systems theory (see e.g. [10, Theorem 2.3.3]), the strong limit
$\partial_{\tau}\psi_{(\tau,t)}:=\lim_{h\to 0}\frac{\psi_{(\tau+h,t)}-\psi_{(\tau,t)}}{h}$ (4.14)
is well-defined in $\pazocal{X}$ for every $(\tau,t)\in[0,T]\times[0,T]$ under our assumptions. Besides, one can easily check that for all $(t,y)\in[0,T]\times\mathbb{R}^{d}$, the map $\tau\in[0,T]\mapsto\psi_{(\tau,t)}\circ\Phi_{(t,\tau)}^{u}(y)$ is constant, so that
$\tfrac{\textnormal{d}}{\textnormal{d}\tau}\Big{(}\psi_{(\tau,t)}\circ\Phi_{(t,\tau)}^{u}(y)\Big{)}$ $=\partial_{\tau}\psi_{(\tau,t)}\circ\Phi_{(t,\tau)}^{u}(y)+\nabla_{x}\psi_{(\tau,t)}\circ\Phi_{(t,\tau)}^{u}(y)\cdot\partial_{\tau}\Phi_{(t,\tau)}^{u}(y)$
$=\Big{(}\partial_{\tau}\psi_{(\tau,t)}+\nabla_{x}\psi_{(\tau,t)}\cdot f_{u(\tau)}\Big{)}\circ\Phi_{(t,\tau)}^{u}(y)$
$=0$
for $\mathscr{L}^{1}$-almost every $\tau\in[0,T]$. Noting that in the previous identity, one may choose the particular point $y=\Phi_{(\tau,t)}^{u}(x)$ for some arbitrary $x\in\mathbb{R}^{d}$, it then follows that $(\tau,x)\in[0,T]\times\mathbb{R}^{d}\mapsto\psi_{(\tau,t)}(x)\in\mathbb{R}^{d}$ is the unique classical solution of the Cauchy problem
[…]
in $[0,T]\times\mathbb{R}^{d}$. At this stage, there remains to notice that the map
$(\tau,x)\in[0,T]\times\mathbb{R}^{d}\mapsto\nabla_{x}\psi_{(\tau,t)}(x)\cdot f_{u(\tau)}(x)\in\mathbb{R}^{d}$
is $\mathscr{L}^{1}$-measurable in $\tau\in[0,T]$ as well as continuous in $x\in\mathbb{R}^{d}$, which implies in particular that its functional lift $\tau\in[0,T]\mapsto\nabla_{x}\psi_{(\tau,t)}\cdot f_{u(\tau)}\in\pazocal{X}$ is $\mathscr{L}^{1}$-measurable, see e.g. [38, Page 511]. The latter fact together with (4.14) and (4.15) allows us to conclude that $\tau\in[0,T]\mapsto\psi_{(\tau,t)}\in\pazocal{D}$ solves (4.13).
To prove the converse implication, observe that one may rewrite the evaluation of the Liouville operator along the curve $\tau\in[0,T]\mapsto\psi_{(\tau,t)}\in\pazocal{D}$ as
$\pazocal{L}(\psi_{(\tau,t)})=\Big{\{}\Psi_{t}(\tau,u)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}$
wherein $\Psi_{t}:(\tau,u)\in[0,T]\times U\mapsto\nabla_{x}\psi_{(\tau,t)}\cdot f_{u}\in\pazocal{X}$. It is clear from Hypothesis (H) that the function $u\in U\mapsto\Psi_{t}(\tau,u)\in\pazocal{X}$ is continuous for $\mathscr{L}^{1}$-almost every $\tau\in[0,T]$, and we just showed $\tau\in[0,T]\mapsto\Psi_{t}(\tau,u)\in\pazocal{X}$ is $\mathscr{L}^{1}$-measurable. Thus, because $\tau\in[0,T]\mapsto\partial_{\tau}\psi_{(\tau,t)}\in\pazocal{X}$ is an $\mathscr{L}^{1}$-measurable map and since the completion of $(\pazocal{X},\parallel\cdot\parallel_{\pazocal{X}})$ is a separable Banach space, if follows from the measurable selection principle of Theorem 2.10 that
$-\partial_{\tau}\psi_{(\tau,t)}=\Psi_{t}(u_{\varphi}(\tau))=\nabla_{x}\psi_{(\tau,t)}\cdot f_{u_{\varphi}(\tau)}$ (4.16)
for $\mathscr{L}^{1}$-almost every $\tau\in[0,T]$ and some admissible control signal $u_{\varphi}(\cdot)\in\pazocal{U}$. Owing to the regularity assumptions posited above, it follows from classical well-posedness results for transport equations (see e.g. [1, Proposition 2.3]) that
$\psi_{(\tau,t)}(x)=\varphi\circ\Phi_{(\tau,t)}^{u_{\varphi}}(x)$
for all $(\tau,x)\in[0,T]\times\mathbb{R}^{d}$, which concludes the proof of our claim. ∎
Remark 4.7 (Comparing Proposition 4.6 with its counterpart in classical Koopman theory).
In the familiar situation in which $(\pazocal{K}_{t}(\varphi))_{t\in[0,T]}$ is the usual Koopman semigroup generated by a single family of flows $(\Phi_{t})_{t\in[0,T]}\subset C^{0}(\mathbb{R}^{d},\mathbb{R}^{d})$, it is well-known (see e.g. [33, Section 7.6]) that the curve of observables $t\in[0,T]\mapsto\psi_{t}:=\pazocal{K}_{t}(\varphi)\in\pazocal{X}$ is the unique strong solution of the Koopman dynamics
[…]
In the previous expression, the right-hand side coincides with the evaluation of the classical Liouville operator $\pazocal{L}:\varphi\in\pazocal{D}\mapsto\nabla_{x}\varphi\cdot f\in\pazocal{X}$ along $(\psi_{t})_{t\in[0,T]}$. It is worth noting that this result is essentially contained in Proposition 4.6. Indeed, in the autonomous case, the Koopman observables take the simpler form $\psi_{(\tau,t)}=\varphi\circ\Phi_{t-\tau}$, and in particular
$\partial_{\tau}\psi_{(\tau,t)}=-\partial_{t}\psi_{(\tau,t)}\in-\pazocal{L}(\psi_{(\tau,t)})=\big{\{}-\nabla_{x}\psi_{(\tau,t)}\cdot f\,\big{\}}.$
Hence, up to redefining the time variable, the set-valued dynamics presented in (4.13) reduces to (4.17).
Remark 4.8 (Pointwise versus functional Koopman dynamics).
In Proposition 4.6, we proved that when the controlled vector fields are sufficiently smooth, the set-valued Liouville operator encodes the dynamics of Koopman observables in a functional sense, as it generates solution curves to a differential inclusion in the space of observables. However, much like what is exposed in [33, Remark 7.6.1], these operators could also be used to describe the pointwise evolution of Koopman observables. Indeed, under Hypotheses (H), if follows from the usual characteristic formula along with Rademacher’s theorem (see e.g. [2, Theorem 2.14]) that $(\tau,x)\in[0,T]\times\mathbb{R}^{d}\mapsto\psi_{(\tau,t)}(x):=\varphi\circ\Phi_{(\tau,t)}^{u}(x)\in\mathbb{R}^{d}$ satisfies
$\partial_{\tau}\psi_{(\tau,t)}(x)+\nabla_{x}\psi_{(\tau,t)}(x)\cdot f_{u(\tau)}(x)=0$
for almost every $\tau,t\in[0,T]$ and $x\in\mathbb{R}^{d}$. This can be informally thought of as a pointwise inclusion of the form $\partial_{\tau}\psi_{(\tau,t)}(x)\in-\pazocal{L}(\psi_{(\tau,t)})(x)$ for almost every $x\in\mathbb{R}^{d}$, and should in all likelihood correspond to a weak version of the functional dynamics (4.13).
4.2 Set-valued Perron-Frobenius operators and adjoint Koopman dynamics
In this section, we propose a formal definition for the set-valued Perron-Frobenius operator, and precisely discuss its relationship with the Koopman semigroup as well as that of its infinitesimal generator with the Liouville operator. In this context, we will frequently resort to the notion of image measure under the action of a Borel map introduced in (2.2) above.
Definition 4.9 (Set-valued Perron-Frobenius operators).
We define the set-valued Perron-Frobenius operators $\pazocal{P}_{(\tau,t)}:\pazocal{X}^{*}\rightrightarrows\pazocal{X}^{*}$ by
$\pazocal{P}_{(\tau,t)}(\mu):=\Big{\{}\Phi^{u}_{(\tau,t)\sharp\,}\mu\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u(\cdot)\in\pazocal{U}\Big{\}}\subset\pazocal{X}^{*}$ (4.18)
for all times $\tau,t\in[0,T]$ and each $\mu\in\pazocal{X}^{*}$.
In what follows, we provide mathematical grounding for this definition by showing that the set-valued Perron-Frobenius operators are related to suitable adjoints of the set-valued Koopman operators, whose definition is inspired by the work [26].
Definition 4.10 (Adjoint of closed processes).
Given two locally convex topological vector spaces $\pazocal{X}$ and $\pazocal{Y}$, we define the adjoint $\pazocal{F}^{*}:\pazocal{Y}^{*}\rightrightarrows\pazocal{X}^{*}$ of a closed process $\pazocal{F}:\pazocal{X}\rightrightarrows\pazocal{Y}$ by
$\pazocal{F}^{*}(\mu):=\bigg{\{}\nu\in\pazocal{X}^{*}\leavevmode\nobreak\ \,\text{s.t.}\leavevmode\nobreak\ \langle\mu,x\rangle_{\pazocal{X}}\leq\sup_{y\in\pazocal{F}(x)}\langle\nu,y\rangle_{\pazocal{Y}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for all $x\in\pazocal{X}$}\bigg{\}}$
for all $\mu\in\pazocal{Y}^{*}$.
Example 4.11 (Adjoints of fans generated by convex families of linear operators).
To see again why the proposed notion of adjoint is natural and compatible with its single-valued counterpart, suppose that $(\pazocal{X},\parallel\cdot\parallel_{\pazocal{X}})$ and $(\pazocal{Y},\parallel\cdot\parallel_{\pazocal{Y}})$ are Banach spaces and let $\pazocal{F}(x)=\{Ax\}$ for each $x\in\pazocal{X}$ and some bounded linear map $A:\pazocal{X}\to\pazocal{Y}$. Then, one may trivially check that
$\sup_{y\in\pazocal{F}(x)}\langle\mu,y\rangle_{\pazocal{Y}}=\langle\mu,Ax\rangle_{\pazocal{Y}}=\langle A^{*}\mu,x\rangle_{\pazocal{X}}$
for each $\mu\in\pazocal{Y}^{*}$ and every $x\in\pazocal{X}^{*}$, and it directly follows from Definition 4.10 that
$\pazocal{F}^{*}(\mu)=\{A^{*}\mu\}.$
More importantly, it is shown in [26] that if $\pazocal{A}$ is a convex and (weakly) closed collection of bounded linear operators from $\pazocal{X}$ into $\pazocal{Y}$, then the adjoint of the set-valued mapping defined by $\pazocal{F}(x):=\{Ax\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ A\in\pazocal{A}\}$ for each $x\in\pazocal{X}$ writes as $\pazocal{F}^{*}(\mu)=\{A^{*}\mu\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ A\in\pazocal{A}\}$ for all $\mu\in\pazocal{Y}^{*}$.
Theorem 4.12 (Characterisation of the adjoints of the Liouville and Koopman operators).
Suppose that Hypotheses (H) hold. Then, one has that
$\pazocal{L}^{*}(\mu)=\overline{\textnormal{co}}\hskip 1.42271pt\Big{\{}-\textnormal{div}_{x}(f_{u}\mu)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}\qquad\text{and}\qquad\pazocal{K}^{*}_{(\tau,t)}(\mu)=\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{P}_{(\tau,t)}(\mu)$ (4.19)
for all $\mu\in\pazocal{X}^{*}$, where the convex hulls are taken with respect to the weak- topology. In particular, if Hypothesis (C) holds, then
$\pazocal{L}^{*}(\mu)=\Big{\{}-\textnormal{div}_{x}(f_{u}\mu)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}$
for all $\mu\in\pazocal{X}^{*}$.
Proof.
The proof of this result relies on a general characterisation of closed convex hulls in locally convex spaces in terms of support functions, based itself on Hahn-Banach’s separation principle. First, we start by noting that given any $\varphi\in\pazocal{D}$ and each $\psi:=\nabla\varphi\cdot f_{u}\in\pazocal{L}(\varphi)$ with $u\in U$, there holds
$\langle\mu,\psi\rangle_{\pazocal{X}}$ $=\langle\mu,\nabla\varphi\cdot f_{u}\rangle_{\pazocal{X}}$
$=\int_{\mathbb{R}^{d}}\nabla\varphi(x)\cdot f_{u}(x)\textnormal{d}\mu(x)=\langle-\textnormal{div}_{x}(f_{u}\mu),\varphi\rangle_{\pazocal{D}}$
for all $\mu\in\pazocal{X}^{*}$, by the definition (2.1) of the divergence distribution. This implies in particular that
$\pazocal{L}^{*}(\mu)$ =$=\bigg{\{}\nu\in\pazocal{D}^{*}\leavevmode\nobreak\ \,\text{s.t.}\leavevmode\nobreak\ \langle\nu,\varphi\rangle_{\pazocal{D}}\leq\sup_{\psi\in\pazocal{L}(\varphi)}\langle\mu,\psi\rangle_{\pazocal{X}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for all $\varphi\in\pazocal{D}$}\bigg{\}}$
=$=\bigg{\{}\nu\in\pazocal{D}^{*}\leavevmode\nobreak\ \,\text{s.t.}\leavevmode\nobreak\ \langle\nu,\varphi\rangle_{\pazocal{D}}\leq\sup_{u\in U}\langle-\textnormal{div}_{x}(f_{u}\mu),\varphi\rangle_{\pazocal{D}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for all $\varphi\in\pazocal{D}$}\bigg{\}}.$
At this stage, recalling that the dual of $\pazocal{D}^{*}$ seen as a topological vector space endowed with the weak- topology is $\pazocal{D}$ itself (see e.g. [11, Proposition 3.14]), it follows from a direct adaptation to locally convex spaces of the characterisation of convex hulls provided in [5, Theorem 2.4.2] that
$\pazocal{L}^{*}(\mu)=\overline{\textnormal{co}}\hskip 1.42271pt\Big{\{}-\textnormal{div}_{x}(f_{u}\mu)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}.$
We now turn our attention to the adjoints of the Koopman semigroup. Similarly to what precedes, given $\varphi\in\pazocal{X}$ and $\psi_{(\tau,t)}=\varphi\circ\Phi_{(\tau,t)}^{u}\in\pazocal{K}_{(\tau,t)}(\varphi)$ with $u(\cdot)\in\pazocal{U}$, we start by observing that
$\langle\mu,\psi_{(\tau,t)}\rangle_{\pazocal{X}}$ $=\int_{\mathbb{R}^{d}}\varphi\circ\Phi_{(\tau,t)}^{u}(x)\textnormal{d}\mu(x)$
$=\int_{\mathbb{R}^{d}}\varphi(x)\textnormal{d}(\Phi_{(\tau,t)\sharp\,}^{u}\mu)(x)=\langle\Phi_{(\tau,t)\sharp\,}^{u}\mu,\varphi\rangle_{\pazocal{X}}$
for all $\mu\in\pazocal{X}^{*}$. Then, following the exact same reasoning as above, one may conclude that
$\pazocal{K}^{*}_{(\tau,t)}(\mu)$ =$=\bigg{\{}\nu_{(\tau,t)}\in\pazocal{X}^{*}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ \langle\nu_{(\tau,t)},\varphi\rangle_{\pazocal{X}}\leq\sup_{\psi_{(\tau,t)}\in\pazocal{K}_{(\tau,t)}(\varphi)}\langle\mu,\psi_{(\tau,t)}\rangle_{\pazocal{X}}\bigg{\}}$
=$=\bigg{\{}\nu_{(\tau,t)}\in\pazocal{X}^{*}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ \langle\nu_{(\tau,t)},\varphi\rangle_{\pazocal{X}}\leq\sup_{u(\cdot)\in\pazocal{U}}\langle\Phi_{(\tau,t)\sharp\,}^{u}\mu,\varphi\rangle_{\pazocal{X}}\bigg{\}}=\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{P}_{(\tau,t)}(\mu)$
by Hahn-Banach’s separation principle applied to $\pazocal{X}^{*}$ endowed with the weak- topology. Lastly, if Hypothesis (C) holds, it may be easily verified that
$\pazocal{L}^{*}(\mu)$ $=\overline{\textnormal{co}}\hskip 1.42271pt\Big{\{}-\textnormal{div}_{x}(f\mu)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ f\in\pazocal{F}\Big{\}}$
[…]
which then concludes the proof. ∎
We end this section by showing that under our working assumptions, the infinitesimal generator of the Perron-Frobenius semigroup [^3] Therein, we make a small abuse of notation by writing Kuratowski-Painlevé limits although the weak- topology is not metrisable. This could be made rigorous by redefining the latter notion in terms of nets and neighbourhoods, but we preferred to avoid unnecessary technicalities and convey more directly the underlying intuition. is the adjoint of the Liouville operator.
Theorem 4.13 (Infinitesimal generator of the Perron-Frobenius semigroup).
Suppose that Hypotheses (H) hold. Then, the following inclusions
$\Big{\{}-\textnormal{div}_{x}(f_{u}\mu)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}\subset\underset{\leavevmode\nobreak\ t\to\tau}{\textnormal{Liminf}}\,\frac{\pazocal{P}_{(\tau,t)}(\mu)-\mu}{t-\tau}$ (4.20)
and
$\underset{\leavevmode\nobreak\ t\to\tau}{\textnormal{Limsup}}\,\frac{\pazocal{P}_{(\tau,t)}(\mu)-\mu}{t}\subset\overline{\textnormal{co}}\hskip 1.42271pt\Big{\{}-\textnormal{div}_{x}(f_{u}\mu)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}$ (4.21)
are satisfied for each $\mu\in\pazocal{X}^{*}$. Furthermore, if Hypothesis (C) holds, then the adjoint of the set-valued Liouville operator is the infinitesimal generator of the Perron-Frobenius semigroup, in the sense that
$\underset{\leavevmode\nobreak\ t\to\tau}{\textnormal{Lim}}\frac{\pazocal{P}_{(\tau,t)}(\mu)-\mu}{t-\tau}=\pazocal{L}^{*}(\mu)$ (4.22)
with the limit being taken in the weak-* topology for all $\mu\in\pazocal{X}^{*}$.
Proof.
Following Theorem 4.12, we start by noting that if Hypothesis (C) holds, then
$\pazocal{L}^{*}(\mu)=\Big{\{}-\textnormal{div}_{x}(f_{u}\mu)\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}}$
and it directly stems from the inclusions (4.20) and (4.21) that (4.22) holds, namely that the adjoint of the Liouville operator is the infinitesimal generator of the Perron-Frobenius operator. We then seek to establish the liminf inclusion (4.20), which amounts to showing that given some $\nu:=-\textnormal{div}_{x}(f_{\bar{u}}\mu)\in\pazocal{D}^{*}$, there exists a control signal $u(\cdot)\in\pazocal{U}$ for which
$\lim_{t\to\tau}\frac{\Phi_{(\tau,t)\sharp\,}^{u}\mu-\mu}{t-\tau}=-\textnormal{div}_{x}(f_{\bar{u}}\mu)$
in the weak- topology. To this end, we simply consider the constant control signal given by $u(t):=\bar{u}$ for all times $t\in[0,T]$, and observe that in this case, it holds for any $\zeta\in\pazocal{D}$ that
$\lim_{t\to\tau}\big{\langle}\tfrac{1}{t-\tau}\big{(}\Phi_{(\tau,t)\sharp\,}^{\bar{u}}\mu-\mu\big{)},\zeta\big{\rangle}_{\pazocal{X}}$ $=\lim_{t\to\tau}\int_{\mathbb{R}^{d}}\tfrac{1}{t-\tau}\Big{(}\zeta\circ\Phi_{(\tau,t)}^{\bar{u}}(x)-\zeta(x)\Big{)}\textnormal{d}\mu(x)$
$=\int_{\mathbb{R}^{d}}\nabla_{x}\zeta(x)\cdot f_{\bar{u}}(x)\,\textnormal{d}\mu(x)$
$=\big{\langle}-\textnormal{div}_{x}(f_{\bar{u}}\,\mu),\zeta\big{\rangle}_{\pazocal{D}}$
by an application of Lebesgue’s dominated convergence theorem, which is licit since $\zeta\in C^{1}_{c}(\mathbb{R}^{d},\mathbb{C})$, thereby yielding the liminf inclusion. To prove the limsup inclusion (4.21), one needs to show
$\lim_{t_{n}\to\tau}\frac{\Phi_{(\tau,t_{n})\sharp\,}^{u_{n}}\mu-\mu}{t_{n}-\tau}\in\overline{\textnormal{co}}\hskip 1.42271pt\pazocal{L}^{*}(\mu)$
for every pair of sequences $(u_{n}(\cdot))\subset\pazocal{U}$ and $(t_{n})\subset[0,T]\setminus\{\tau\}$ along which this limit exists in the weak- topology. Given some $\zeta\in\pazocal{D}$, we consider the compact set
$K_{\zeta}:=\overline{\bigcup_{u(\cdot)\in\pazocal{U}}\bigcup_{\tau,t\in[0,T]}\Phi^{u}_{(t,\tau)}(\textnormal{supp}(\zeta))}$
and note by reproducing the reasoning detailed in the proof of Theorem 4.3 that there exists some $\bar{u}\in U$ such that
$\sup_{x\in K_{\zeta}}\big{|}\Phi^{u_{n}}_{(\tau,t_{n})}(x)-x-(t_{n}-\tau)f_{\bar{u}}(x)\big{|}\leq o(|t_{n}-\tau|)$
along a subsequence $t_{n}\to\tau$ that we do not relabel. The latter identity further implies that
$\lim_{t_{n}\to\tau}\big{\langle}\tfrac{1}{t_{n}-\tau}\big{(}\Phi^{u_{n}}_{(\tau,t_{n})\sharp\,}\mu-\mu\big{)},\zeta\big{\rangle}_{\pazocal{X}}$ $=\lim_{t_{n}\to\tau}\int_{\mathbb{R}^{d}}\tfrac{1}{t_{n}-\tau}\Big{(}\zeta\circ\Phi^{u}_{(\tau,t_{n})}(x)-\zeta(x)\Big{)}\textnormal{d}\mu(x)$
$=\int_{\mathbb{R}^{d}}\nabla_{x}\zeta(x)\cdot f_{\bar{u}}(x)\textnormal{d}\mu(x)$
$=\big{\langle}-\textnormal{div}_{x}(f_{\bar{u}}\,\mu),\zeta\big{\rangle}_{\pazocal{D}}$
by Lebesgue’s dominated convergence theorem along with the definition (2.1) of the divergence distribution and the fact that $\zeta\in C^{1}_{c}(\mathbb{R}^{d},\mathbb{C})$, which concludes the proof of our claim. ∎
As a consequence of Theorem 4.13, it can be shown [^4] The precise argument involves abstract measurable selection theorems adapted to weak- topologies – which are not metrisable but Souslin [47, Theorem 7 page 112] – that may be found e.g. in [13], and exceeds the scope of this article. by adapting the arguments of Proposition 4.6 that a curve $t\in[0,T]\mapsto\mu_{(\tau,t)}\in\pazocal{P}_{(\tau,t)}(\mu)$ is of the form $\mu_{(\tau,t)}=\Phi_{(\tau,t)\sharp\,}^{u}\mu$ for some fixed control signal $u(\cdot)\in\pazocal{U}$ if and only if it is a weak- solution of the differential inclusion
$&\partial_{t}\mu_{(\tau,t)}\in\Big{\{}-\textnormal{div}_{x}(f_{u}\mu_{(\tau,t)})\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ u\in U\Big{\}},\\ &\mu_{(\tau,\tau)}=\mu.\end{aligned}\right.$
Interestingly, it follows from the characterisation derived in Theorem 4.12 that, when the velocities are convex, the adjoint Koopman dynamics is essentially a differential inclusion in the space of measures as introduced by the first author in [8, 9] to investigate mean-field control problems.
4.3 Point spectra of the set-valued Koopman and Liouville operators
In this section, we briefly study the interplay between the point spectrum of the set-valued Koopman and Liouville operators, understood in the following sense.
Definition 4.14 (Eigenvalues and eigenvectors of closed processes).
Given two Banach spaces $(\pazocal{X},\parallel\cdot\parallel_{\pazocal{X}})$ and $(\pazocal{Y},\parallel\cdot\parallel_{\pazocal{Y}})$, we say that a complex number $\lambda\in\mathbb{C}$ is an eigenvalue of a closed process $\pazocal{F}:\pazocal{X}\rightrightarrows\pazocal{Y}$ if there exists an eigenvector $x_{\lambda}\in\pazocal{X}\setminus\{0\}$ such that
$\lambda x_{\lambda}\in\pazocal{F}(x_{\lambda}).$
The set of all such eigenvalues is called the point spectrum of $\pazocal{F}$, and is denoted by $\sigma_{p}(\pazocal{F})\subset\mathbb{C}$.
Example 4.15 (Link with the usual notion of point spectrum).
Again, one may easily check that the definition of spectrum introduced above reduces to the usual one when $\pazocal{F}(x)=\{Ax\}$ with $A:\pazocal{X}\to\pazocal{Y}$ linear and bounded, since then $\lambda x_{\lambda}\in\pazocal{F}(x_{\lambda})=\{Ax_{\lambda}\}$ if and only if $Ax_{\lambda}=\lambda x_{\lambda}$ for each $\lambda\in\sigma_{p}(\pazocal{F})$.
Below, we prove a set-valued version of the classical spectral mapping theorem (see e.g. [19, Chapter IV - Theorem 3.7]), which relates the point spectra of the Liouville and Koopman operators.
Theorem 4.16 (Relationship between the eigenvalues of the Koopman and Liouville operators).
Suppose that Hypotheses (H) hold and fix some $\tau\in[0,T]$. Then for all times $t\in[0,T]$, it holds that
$e^{\sigma_{p}(\pazocal{L})(t-\tau)}\subset\sigma_{p}(\pazocal{K}_{(\tau,t)}),$ (4.23)
namely if an observable $\varphi_{\lambda}\in\pazocal{D}$ is an eigenfunction of $\pazocal{L}:\pazocal{D}\rightrightarrows\pazocal{X}$ with eigenvalue $\lambda\in\sigma_{p}(\pazocal{L})$, then it is an eigenfunction of $\pazocal{K}_{(\tau,t)}:\pazocal{X}\rightrightarrows\pazocal{X}$ with eigenvalue $e^{\lambda(t-\tau)}\in\sigma_{p}(\pazocal{K}_{(\tau,t)})$. Furthermore, if Hypothesis (C) holds, then for each measurable curve $t\in[0,T]\mapsto\lambda_{(\tau,t)}\in\mathbb{C}$ satisfying
$\lambda_{(\tau,t)}\varphi_{\lambda}\in\pazocal{K}_{(\tau,t)}(\varphi_{\lambda})$
for some given time-independent $\varphi_{\lambda}\in\pazocal{D}$, there exists a sequence $t_{n}\to\tau$ for which
$\lim_{n\to+\infty}\lambda_{(\tau,t_{n})}=1\qquad\text{and}\qquad\lim_{n\to+\infty}\frac{\lambda_{(\tau,t_{n})}-1}{t_{n}-\tau}\in\sigma_{p}(\pazocal{L}).$ (4.24)
In that case, the converse spectral inclusion
$\bigg{\{}\lambda_{(\tau,t)}\in\mathbb{C}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ \lambda_{(\tau,t)}\varphi_{\lambda}\in\pazocal{K}_{(\tau,t)}(\varphi_{\lambda})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \lambda_{(\tau,t)}=e^{\lambda(t-\tau)}\leavevmode\nobreak\ \text{for some}\leavevmode\nobreak\ \lambda\in\mathbb{C}\bigg{\}}\subset e^{\sigma_{p}(\pazocal{L})(t-\tau)}$ (4.25)
holds for all times $\tau,t\in[0,T]$.
Proof.
We start by proving the first inclusion (4.23). By Theorem 4.3, the fact that $\lambda\varphi_{\lambda}\in\pazocal{L}(\varphi_{\lambda})$ for some $\lambda\in\sigma_{p}(\pazocal{L})$ and $\varphi_{\lambda}\in\pazocal{D}$ is tantamount to the existence of some control value $\bar{u}_{\lambda}\in U$ for which
$\lambda\varphi_{\lambda}=\nabla_{x}\varphi_{\lambda}\cdot f_{\bar{u}_{\lambda}}.$ (4.26)
Hence, considering the constant control signal defined by $\bar{u}_{\lambda}(t):=\bar{u}_{\lambda}$ for all times $t\in[0,T]$, one has
$\tfrac{\textnormal{d}}{\textnormal{d}t}{}\Big{(}\varphi\circ\Phi^{u_{\lambda}}_{(\tau,t)}(x)\Big{)}=\langle\nabla_{x}\varphi_{\lambda},f_{\bar{u}_{\lambda}}\rangle\circ\Phi_{(\tau,t)}^{u_{\lambda}}(x)=\lambda\varphi_{\lambda}\circ\Phi_{(\tau,t)}^{u_{\lambda}}(x)$
for each $x\in\mathbb{R}^{d}$, and it follows from a standard Cauchy-Lipschitz uniqueness argument that
$\varphi\circ\Phi^{u_{\lambda}}_{(\tau,t)}(x)=e^{\lambda(t-\tau)}\varphi_{\lambda}(x)$
for all $(t,x)\in Z[0,T]\times\mathbb{R}^{d}$. The thesis follows then from the observation that $\varphi\circ\Phi^{u_{\lambda}}_{(\tau,t)}\in\pazocal{K}_{(\tau,t)}(\varphi_{\lambda})$, by the very definition of the set-valued Koopman operators.
Conversely, let us assume that there exist $t\mapsto\lambda_{(\tau,t)}\in\mathbb{C}$ and $\varphi_{\lambda}\in\pazocal{D}$ such that $\lambda_{(\tau,t)}\varphi_{\lambda}\in\pazocal{K}_{(\tau,t)}(\varphi_{\lambda})$ for $\mathscr{L}^{1}$-almost every $t\in[0,T]$, and suppose that Hypothesis (C) holds. By Proposition 3.5, there exists an $\mathscr{L}^{1}\times\mathscr{L}^{1}$-measurable map $(\tau,t)\in[0,T]\times[0,T]\mapsto u_{(\tau,t)}^{\lambda}(\cdot)\in\pazocal{U}$ such that
$\lambda_{(\tau,t)}\varphi_{\lambda}=\varphi_{\lambda}\circ\Phi_{(\tau,t)}^{u_{(\tau,t)}^{\lambda}}\in\pazocal{K}_{(\tau,t)}(\varphi_{\lambda})$ (4.27)
for $\mathscr{L}^{1}$-almost every $t\in[0,T]$. By repeating the arguments yielding the limsup inclusion of Theorem 4.3, one may infer the existence of an element $\bar{u}_{\lambda}\subset U$ along with a sequence $t_{n}\to\tau^{+}$ such that
$\varphi_{\lambda}\circ\Phi_{(\tau,t_{n})}^{u_{(\tau,t_{n})}^{\lambda}}=\varphi_{\lambda}+(t_{n}-\tau)\nabla_{x}\varphi_{\lambda}\cdot f_{\bar{u}_{\lambda}}+o(|t_{n}-\tau|).$
This along with (4.27) implies that $\lim_{n\to+\infty}\lambda_{(\tau,t_{n})}=1$, as well as
$\lim_{n\to+\infty}\frac{\lambda_{(\tau,t_{n})}-1}{t_{n}-\tau}\varphi_{\lambda}$ $=\lim_{n\to+\infty}\frac{\varphi_{\lambda}\circ\Phi_{(\tau,t_{n})}^{u_{(\tau,t_{n})}^{\lambda}}-\varphi_{\lambda}}{t_{n}-\tau}$
$=\nabla_{x}\varphi_{\lambda}\cdot f_{\bar{u}_{\lambda}},$
which belongs to $\pazocal{L}(\varphi_{\lambda})$ as a consequence of Theorem 4.3. Thus, we have shown the limit inclusion (4.24), from which the spectral inclusion (4.25) easily follows. ∎
Example 4.17 (The spectral mapping theorem for linear feedback controls).
To illustrate the previous theorem, let us focus on the simple example of a linear time-invariant system with feedback controls
$\dot{x}(t)=(A+BK)x(t)$
wherein $(A,B)\in\mathbb{R}^{d\times d}\times\mathbb{R}^{d\times m}$ and $K\in\pazocal{K}_{\textnormal{adm}}$ with $\pazocal{K}_{\textnormal{adm}}\subset\mathbb{R}^{m\times d}$ being a compact set. In practice, one may consider for instance that $\pazocal{K}_{\textnormal{adm}}:=\{K_{1},\dots,K_{n}\}$ is comprised of finitely many feedback matrices, or maybe fix $\pazocal{K}_{\textnormal{adm}}:=\big{\{}K\in\mathbb{R}^{m\times d}\leavevmode\nobreak\ \,\textnormal{s.t.}\;\parallel K\parallel_{\mathbb{F}}\,\leq 1\big{\}}$ as the closed unit ball for the Frobenius norm. In this context, the set-valued Koopman and Liouville operators can be computed explicitly as
$\pazocal{K}_{(\tau,t)}:\varphi\in\pazocal{X}\rightrightarrows\bigg{\{}x\in\mathbb{R}^{d}\mapsto\varphi\bigg{(}\exp\Big{(}(t-\tau)(A+BK)\Big{)}x\bigg{)}\in\mathbb{C}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ K\in\pazocal{K}_{\textnormal{adm}}\bigg{\}}\subset\pazocal{X}$
and
$\pazocal{L}:\varphi\in\pazocal{D}\rightrightarrows\bigg{\{}x\in\mathbb{R}^{d}\mapsto\nabla_{x}\varphi(x)\cdot\Big{(}(A+BK)x\Big{)}\in\mathbb{C}\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ K\in\pazocal{K}_{\textnormal{adm}}\bigg{\}}\subset\pazocal{X}.$
In this context, a map $\varphi_{\lambda}\in\pazocal{D}$ is an eigenvalue of the Liouville operator if there exists some $\lambda\in\mathbb{C}$ along with a $K_{\lambda}\in\pazocal{K}_{\textnormal{adm}}$ such that
$\lambda\varphi_{\lambda}(x)=\nabla_{x}\varphi_{\lambda}(x)\cdot(A+BK_{\lambda})x$
for all $x\in\mathbb{R}^{d}$. Following [35], one may for instance consider candidates linear eigenfunctions of the form $\varphi_{\lambda}(x):=e_{\lambda}\cdot x$, with $\lambda\in\mathbb{C}$ and $e_{\lambda}\in\mathbb{C}^{d}$ being respectively an eigenvalue and eigenvector of $(A+BK_{\lambda})^{*}\in\mathbb{C}^{d\times d}$. One may then check that $e^{\lambda t}\in\sigma_{p}(\pazocal{K}_{(0,t)})$ for each $\lambda\in\sigma_{p}(\pazocal{L})$, and vice versa.
We end this section by a simple proposition which shows how one may produce new eigenfunctions and eigenvalues of the set-valued Liouville operator by combining already known ones.
Proposition 4.18 (On the structure of the point spectrum).
Let $\varphi_{\lambda_{1}},\varphi_{\lambda_{2}}\in\pazocal{D}$ be two eigenfunctions of $\pazocal{L}:\pazocal{D}\rightrightarrows\pazocal{X}$ associated with the same control value $u_{\lambda}\in U$, namely
$\lambda_{1}\varphi_{\lambda_{1}}=\nabla_{x}\varphi_{\lambda_{1}}\cdot f_{u_{\lambda}}\qquad\text{and}\qquad\lambda_{2}\varphi_{\lambda_{2}}=\nabla_{x}\varphi_{\lambda_{2}}\cdot f_{u_{\lambda}}.$
Then for every $\alpha_{1},\alpha_{2}\in\mathbb{R}$ such that $\varphi_{\lambda_{1}}^{\alpha_{1}}\varphi_{\lambda_{2}}^{\alpha_{2}}\in\pazocal{D}$, the latter is an eigenfunction of $\pazocal{L}:\pazocal{D}\rightrightarrows\pazocal{X}$ with the eigenvalue $\alpha_{1}\lambda_{1}+\alpha_{2}\lambda_{2}\in\sigma_{p}(\pazocal{L})$.
Proof.
Given a pair of real numbers $\alpha_{1},\alpha_{2}\in\mathbb{R}$ for which $\varphi_{\lambda_{1}}^{\alpha_{1}}\varphi_{\lambda_{2}}^{\alpha_{2}}\in\pazocal{D}$, one easily gets that
$\nabla_{x}(\varphi_{\lambda_{1}}^{\alpha_{1}}\varphi_{\lambda_{2}}^{\alpha_{2}})\cdot f_{u_{\lambda}}$ $=\alpha_{1}\varphi_{\lambda_{1}}^{\alpha_{1}-1}\varphi_{\lambda_{2}}^{\alpha_{2}}\;\nabla_{x}\varphi_{\lambda_{1}}\cdot f_{u_{\lambda}}+\alpha_{2}\varphi_{\lambda_{1}}^{\alpha_{1}}\varphi_{\lambda_{2}}^{\alpha_{2}-1}\;\nabla_{x}\varphi_{\lambda_{2}}\cdot f_{u_{\lambda}}$
$=(\alpha_{1}\lambda_{1}+\alpha_{2}\lambda_{2})\varphi_{\lambda_{1}}^{\alpha_{1}}\varphi_{\lambda_{2}}^{\alpha_{2}},$
which proves the statement. ∎
5 Conclusion
In this paper, we proposed a new mathematical framework for Koopman operators associated with control systems, formulated using the tools and concepts of set-valued analysis. The main rationale behind this work was to provide a sound Koopman theory for systems with inputs, without assuming that the latter be subjected to some dynamical evolution. The next natural step in this line of research is to develop numerical methods based on the theoretical foundation laid out in this work, thereby giving rise, for example, to a set-valued extension of the classical Dynamic Mode Decomposition.
Appendices
A Proof of Corollary 2.17
In this appendix section, we detail the proof of Corollary 2.17 for the sake of self-containedness.
Proof of Corollary 2.17.
Consider an element $u(\cdot)\in\pazocal{U}$ as well as a sequence $(u_{n}(\cdot))\subset\pazocal{U}$ such that
$\int_{0}^{T}\textnormal{{d}}_{U}(u_{n}(t),u(t))\textnormal{d}t\leavevmode\nobreak\ \underset{n\to+\infty}{\longrightarrow}\leavevmode\nobreak\ 0.$
It then follows from classical results in measure theory (see e.g. [2, Remark 1.18]) that $(u_{n}(\cdot))$ converges to $u(\cdot)$ in measures, that is if we consider the set $\pazocal{A}_{n}^{\delta}:=\big{\{}t\in[0,T]\leavevmode\nobreak\ \,\textnormal{s.t.}\leavevmode\nobreak\ \textnormal{{d}}_{U}(u_{n}(t),u(t))\geq\delta\big{\}}$, it holds
[…]
for each $\delta>0$. In turn, given a compact set $K\subset\mathbb{R}^{d}$, it follows from (2.8) in Theorem 2.16 that
$\sup_{n\geq 1}\big{|}\Phi_{(\tau,t)}^{u_{n}}(x)\big{|}\leq R_{K}$
for all $(\tau,t,x)\in[0,T]\times[0,T]\times K$ and some constant $R_{K}>0$ which only depends on the magnitudes of $m,T$ and $\sup_{x\in K}|x|$. Then, introducing the notation $K^{\prime}:=B(0,R_{K})$, one can use Hypothesis (H)-$(ii)$ to estimate the discrepancy between the flow maps generated by $u_{n}(\cdot)$ and $u(\cdot)$ as
$\big{|}\Phi_{(\tau,t)}^{u}(x)-\Phi_{(\tau,t)}^{u_{n}}(x)\big{|}$ $\leq\int_{\tau}^{t}\Big{|}f\Big{(}\Phi_{(\tau,s)}^{u}(x),u(s)\Big{)}-f\Big{(}\Phi_{(\tau,s)}^{u_{n}}(x),u_{n}(s)\Big{)}\Big{|}\textnormal{d}s$
$\leq\int_{\tau}^{t}\bigg{(}\ell_{K^{\prime}}\big{|}\Phi_{(\tau,s)}^{u}(x)-\Phi_{(\tau,s)}^{u_{n}}(x)\big{|}+\Big{|}f\Big{(}\Phi_{(\tau,s)}^{u}(x),u(s)\Big{)}-f\Big{(}\Phi_{(\tau,s)}^{u}(x),u_{n}(s)\Big{)}\Big{|}\bigg{)}\textnormal{d}s.$
Thence, it directly follows from an application of Grönwall’s lemma that
$\big{|}\Phi_{(\tau,t)}^{u}(x)-\Phi_{(\tau,t)}^{u_{n}}(x)\big{|}\leq\bigg{(}\sup_{x\in K}\int_{0}^{T}\Big{|}f\Big{(}\Phi_{(\tau,s)}^{u}(x),u(s)\Big{)}-f\Big{(}\Phi_{(\tau,s)}^{u}(x),u_{n}(s)\Big{)}\Big{|}\textnormal{d}s\bigg{)}e^{\ell_{K^{\prime}}T}$ (A.2)
for all $(\tau,t,x)\in[0,T]\times[0,T]\times K$ and each $n\geq 1$.
At this stage, given some arbitrary $\varepsilon>0$, it follows from Hypotheses (H) along with the fact that $[0,T]\times[0,T]\times K\times U$ is a compact metric space that there exists some $\delta>0$ for which
$\sup_{(\tau,t,x)\in[0,T]\times[0,T]\times K}\Big{|}f\Big{(}\Phi_{(\tau,t)}^{u}(x),u_{1}\Big{)}-f\Big{(}\Phi_{(\tau,t)}^{u}(x),u_{2}\Big{)}\Big{|}<\frac{\varepsilon}{2Te^{\ell_{K^{\prime}}T}}$ (A.3)
whenever $u_{1},u_{2}\in U$ are such that $\textnormal{{d}}_{U}(u_{1},u_{2})<\delta$. Moreover, following (A.1), there exists for that same $\delta>0$ some integer $N_{\varepsilon,K}\geq 1$ such that
$\mathscr{L}^{1}(\pazocal{A}_{n}^{\delta})<\frac{\varepsilon}{4m(1+R_{K})e^{\ell_{K^{\prime}}T}}$ (A.4)
for each $n\geq N_{\varepsilon}$. Therefore, by combining the information from (A.3) and (A.4), one may infer that
$\int_{0}^{T}\Big{|}f\Big{(}\Phi_{(\tau,s)}^{u}(x),u(s)\Big{)}-f\Big{(}\Phi_{(\tau,s)}^{u}(x),u_{n}(s)\Big{)}\Big{|}\textnormal{d}s$
$\hskip 28.45274pt\leq\int_{\pazocal{A}_{n}^{\delta}}\Big{|}f\Big{(}\Phi_{(\tau,s)}^{u}(x),u(s)\Big{)}-f\Big{(}\Phi_{(\tau,s)}^{u}(x),u_{n}(s)\Big{)}\Big{|}\textnormal{d}s$
$\hskip 41.25641pt+\int_{[0,T]\backslash\pazocal{A}_{n}^{\delta}}\Big{|}f\Big{(}\Phi_{(\tau,s)}^{u}(x),u(s)\Big{)}-f\Big{(}\Phi_{(\tau,s)}^{u}(x),u_{n}(s)\Big{)}\Big{|}\textnormal{d}s$
$\hskip 28.45274pt\leq 2m(1+R_{K})\mathscr{L}^{1}(\pazocal{A}_{n}^{\delta})$
$\hskip 41.25641pt+\sup_{(\tau,s,x)\in[0,T]\times([0,T]\setminus\pazocal{A}_{n}^{\delta})\times K}\Big{|}f\Big{(}\Phi_{(\tau,s)}^{u}(x),u(s)\Big{)}-f\Big{(}\Phi_{(\tau,s)}^{u}(x),u_{n}(s)\Big{)}\Big{|}\,\mathscr{L}^{1}([0,T]\setminus\pazocal{A}_{n}^{\delta})$
$\hskip 28.45274pt\leq\frac{\varepsilon}{e^{\ell_{K^{\prime}}T}}$
for all $x\in K$ and each $n\geq N_{\varepsilon,K}$, where we used Hypothesis (H)-$(i)$. Plugging this last inequality in (A.2) finally yields the uniform estimate
$\sup_{(\tau,t,x)\in[0,T]\times[0,T]\times K}\big{|}\Phi_{(\tau,t)}^{u}(x)-\Phi_{(\tau,t)}^{u_{n}}(x)\big{|}\leq\varepsilon$
valid for each $n\geq N_{\varepsilon,K}$, which concludes the proof since both $K\subset\mathbb{R}^{d}$ and $\varepsilon>0$ are arbitrary. ∎
Acknowledgements
This research was part of the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. This work was also co-funded by the European Union under the project ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590). The authors thank the anonymous referees for their careful reading and comments which greatly contributed to improve the manuscript.